### v1

# When Reasoning Meets Compression: Benchmarking Compressed Large Reasoning Models on Complex Reasoning Tasks

Abstract
Recent open-source large reasoning models (LRMs) exhibit strong performance on complex reasoning tasks, but their large parameter count makes them prohibitively expensive for individuals. The compression of large language models (LLMs) offers an effective solution to reduce cost of computational resources. However, systematic studies on the performance of compressed LLMs in complex reasoning tasks, especially for LRMs, are lacking. Most works on quantization and pruning focus on preserving language modeling performance, while existing distillation works do not comprehensively benchmark student models based on reasoning difficulty or compression impact on knowledge and reasoning. In this paper, we benchmark compressed DeepSeek-R1 models on four different reasoning datasets (AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue), ranging from mathematical to multihop reasoning, using quantization, distillation, and pruning methods. We benchmark 2.51-, 1.73-, and 1.58-bit R1 models that adopt dynamic quantization. We also benchmark distilled R1 models that are based on LLaMA or Qwen and run SparseGPT on them to obtain various sparsity levels. Studying the performance and behavior of compressed LRMs, we report their performance scores and test-time compute (number of tokens spent on each question). Notably, using MuSiQue, we find that parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability, which can inform the choice of compression techniques. Through our empirical analysis of test-time compute, we find that shorter model outputs generally achieve better performance than longer ones across several benchmarks for both R1 and its compressed variants, highlighting the need for more concise reasoning chains.
1 Introduction
Large reasoning models (LRMs) excel at complex reasoning tasks. DeepSeek-R1 (Guo et al., 2025) is the first open-source LRM to successfully replicate OpenAI’s O1-level performance. Trained based on DeepSeek-V3 (DeepSeek-AI et al., 2025), R1 adopts large-scale reinforcement learning without supervised fine-tuning (SFT). Its open-source nature facilitates research on reasoning models (Srivastava et al., 2025; Mondillo et al., 2025). However, due to its large size (671B total parameters), deploying it locally can be costly and even infeasible for individuals, which hinders AI democratization.
Compression of large language models (LLMs) includes quantization, distillation, and pruning. Its goal is to reduce computational resources (e.g., GPU memory and disk space) and provide inference speedup. As the weights of vanilla LLMs are typically in 16-bit floating values (e.g.., FP16 or BF16), quantization means to convert these high-precision values to lower-precision ones. Representative techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization (Lin et al., 2024), and post-training quantization (Frantar et al., 2022). Distillation aims to transfer knowledge from a larger teacher model to a smaller student model under black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings. As for pruning, it aims to sparsify LLMs by introducing 0s to replace nonzero weights. Representative techniques include unstructured pruning (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
Although existing quantization and pruning methods claim to preserve the performance of non-reasoning LLMs after compression such as LLaMA (Touvron et al., 2023) and OPT (Zhang et al., 2022)), exploring these methods on LRMs offers a timely research of democratizing the latest LRMs. Moreover, since these methods mainly adopt perplexity and simple end-task scores for evaluation, the assessment of compressed LRMs will be fundamentally different. Without focusing on basic linguistic capabilities (e.g., simple language understanding and generation), evaluation of compressed reasoning models needs to align with their uncompressed counterparts. Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider compression impact on knowledge and reasoning. A comprehensive evaluation that includes all popular compression methods is valuable.
Therefore, due to the lack of compression research on LRMs, we benchmark the performance and behavior of compressed DeepSeek-R1 on various reasoning tasks in this paper. Our benchmarking framework is shown in Figure 1. We test dynamic quantization (Daniel Han & team, 2023), distillation with SFT (Guo et al., 2025), and SpraseGPT (Frantar & Alistarh, 2023) on R1 (or distilled R1) due to their popularity and effectiveness. Specifically, our chosen dynamic quantization reduces the MoE (Mixture of Experts) layers of R1 model to 2.51-bit, 1.73-bit, and 1.58-bit precision. Testing distillation with SFT, we evaluate the distilled R1 models that adopt LLaMA-70B (Meta AI, ) and Qwen-32B (Qwen et al., 2025) as students. Using SparseGPT, we prune the distilled R1 models to different sparsity levels. We select four complex reasoning benchmarks: mathematical, logical, temporal, and multihop reasoning.
Figure 1: An overview of selected compression methods and tasks. We evaluate R1 and its compressed variants based on their performance scores and test-time compute.
In addition to assessing model performance on selected benchmarks, we study the behavior of compressed R1 models by measuring their test-time compute (number of tokens spent on each question). Specifically, we compare their performance on the shortest and longest outputs. Our major findings are:
  • Parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability. Therefore, both distillation and pruning are discouraged when the end task relies on models’ parametric knowledge. Otherwise, larger student models or low sparsity levels are recommended.
  • In almost all cases, both R1 and its compressed variants achieve higher scores when they spend less compute during inference time. Shorter outputs tend to be more accurate than longer ones across our benchmarks. This highlights the need to reduce verbosity to improve reasoning performance.
2 Related Work
2.1 Quantization
Quantization reduces the number of bits used to represent LLM weights, thereby lowering their precision (Srivastava et al., 2025). Recent survey (Zhu et al., 2024) categorizes quantization methodologies into quantization-aware training (QAT) and post-training quantization (PTQ). QAT requires retraining of model weights to recover performance loss during quantization while PTQ does not require retraining. Recent QAT includes LLM-QAT (Liu et al., 2024a) that adopts distillation to train a quantized LLM, BitDistiller (Du et al., 2024) that develops a self-distillation approach for the full-precision model to act as the teacher of its low-bit counterpart, BitNet (Wang et al., 2023) that proposes a 1-bit Transformer architecture for training LLMs from scratch, and OneBit (Xu et al., 2024) that quantizes LLM weight matrices to 1-bit from a knowledge transfer perspective.
PTQ is more popular in terms of the number of recent publications, because there is no retraining involved. For example, GPTQ (Frantar et al., 2022) is a one-shot weight quantization method that uses approximate second-order information, while AWQ (Lin et al., 2024) leverages activation distribution for finding the salient weight channels to skip. Other PTQ methods include weight-activation quantization (Shao et al., 2024; Yao et al., 2022; Liu et al., 2023) and KV cache quantization (Hooper et al., 2024; Liu et al., 2024b).
2.2 Distillation
Distillation involves two settings: black-box and white-box settings. For black-box setting, teacher model is typically a closed-source LLM and only the outputs of teacher are available for student model. For white-box setting, both weights and output distribution of the teacher model are available. Existing black-box distillation (Huang et al., 2024; Li et al., 2024b; Ho et al., 2023; Huang et al., 2022; Li et al., 2024a) prompts the teacher model to generate a training dataset for the student to learn. Specifically, researchers have started to distill OpenAI’s O1 model (Huang et al., 2024), which marks the beginning of LRMs compression. White-box distillation allows the student model to learn from teacher’s knowledge representation. Works has been done to align the output distribution (Agarwal et al., 2024; Gu et al., 2024) or the hidden representation (Liang et al., 2023) between teacher and student models.
2.3 Pruning
There are unstructured and structured pruning. For unstructured pruning, individual weights are targeted, which leads to irregular sparsity structure. In contrast, structured pruning involves removing entire network components such as channels or layers (Zhang et al., 2024a). Unstructured pruning usually has better compression performance than structured pruning, while it is easier to achieve inference speedup via structured methods (Zhu et al., 2024). Recent unstructured pruning includes one-shot pruning (Frantar & Alistarh, 2023; Sun et al., 2023), global pruning that makes pruning decisions based on all layers (Bai et al., 2024), and domain-specific pruning (Zhang et al., 2024a). Structured pruning includes gradient-based  (Xia et al., 2024; Ma et al., 2023) and non-gradient-based (Ashkboos et al., 2024) methods.
2.4 LRMs
Trained with reinforcement learning, LRMs extends LLMs with advanced reasoning mechanisms (Besta et al., 2025). Popular closed-source LRMs are OpenAI’s o1-mini, o1 (OpenAI et al., 2024), and o3-mini. Open-source LRMs include DeepSeek-R1 and QwQ-32B-Preview (Team, 2024). Since quantization, white-box distillation, and pruning methods require access to model weights, they are not suitable for closed-source LRMs. Only black-box distillation will work on closed-source models.
3 Background
Scope We benchmark quantization, distillation, and pruning methodologies, because we consider them as major LLMs compression paradigms in this paper. Due to lower popularity, other methodologies that can potentially facilitate compression such as low-rank factorization (Sharma et al., 2024; Srebro & Jaakkola, 2003) are not elaborated. When selecting models for compression, we focus exclusively on LRMs.
Recent efforts on LRMs compression As reviewed in Section 2, compression on LRMs (not LLMs) is relatively underexplored. Recently, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights. By combining four ideas (Yu et al., 2024; Dettmers et al., 2023; Ma et al., 2024; Gerganov, 2023), Unsloth quantizes DeepSeek-R1 into four variants: 2.51-, 2.22-, 1.73-, and 1.58-bit models. Here the number of bits (e.g., 2.51-bit) represents the precision of MoE weights, which make up 88% of the total R1 weights. DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation.
Current bottlenecks Very few quantization or pruning works have demonstrated effectiveness on LRMs, as LRMs have only recently emerged. Moreover, evaluation of quantized or pruned LRMs will be different from quantized or pruned LLMs. Current works evaluate quantization and pruning performance primarily using perplexity and simple end tasks, such as the EleutherAI evaluation harness (Gao et al., 2024) and commonsense reasoning. However, compressed LRMs should be assessed on more complex reasoning tasks with varying difficulty levels. As for distillation, although recent distillation works tend to test on more challenging reasoning tasks (compared to other compression literature) such as GSM8K (Cobbe et al., 2021), it is unclear how the compression of LRMs affects models’ parametric knowledge and reasoning capability. Some of them do not comprehensively select diverse reasoning benchmarks. Our benchmarking framework aims to address these bottlenecks.
#Questions #Pass Answer Type Metric Knowledge required?
AIME 2024 30 1 Integer Accuracy False
FOLIO 203 1 True/False/Uncertain Accuracy False
Temporal 250 1 (A)/(B)/(C)/(D) Accuracy False
MuSiQue 100 1 A few words (EM, F1) True
Table 1: Statistics of selected benchmarks: AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue. We use exact match (EM) and F1 to measure model performance on MuSiQue.
4 Experiment Setup
4.1 Reasoning Benchmarks
To address the evaluation bottleneck discussed in Section 3, we select four different reasoning benchmarks with varying levels of difficulty: AIME 2024 (Mathematical Association of America, ) for mathematical reasoning, FOLIO (Han et al., 2024) for logical reasoning, Temporal Sequences of BIG-Bench Hard (Suzgun et al., 2022) for temporal reasoning, and MuSiQue (Trivedi et al., 2022) for multihop reasoning. Table 1 summarizes their statistics. For each model, we do a single pass on every benchmark.
Unlike other benchmarks where we select all questions, we randomly sample 100 out of 1000 from MuSiQue, as it is much larger. This pool of 1000 questions is the same as SiReRAG (Zhang et al., 2025) and HippoRAG (Jimenez Gutierrez et al., 2024). Since MuSiQue requires knowledge memorization besides multihop reasoning, it is popular under retrieval-augmented generation (RAG) setup. Instead of RAG, we follow a closed-book setting (directly prompting LRMs to get final answers) to test models’ parametric knowledge. In this case, unlike other benchmarks that test only reasoning capability, we use MuSiQue to evaluate both reasoning and knowledge memorization. Additional details of selected benchmarks are specified in Appendix A.
4.2 Selected Compression Methods
We decide to compress DeepSeek-R1 due to it large parameter count. For quantization, we select 2.51-, 1.73-, and 1.58-bit models [^1] https://huggingface.co/unsloth/DeepSeek-R1-GGUF by Unsloth. These dynamically quantized models are among the most widely used quantization methods on R1. As for distillation, we select the two largest distilled models that accompany R1: DeepSeek-R1-Distill-Llama-70B [^2] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-32B [^3] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-32B. Since we do not find a pruning method designed for models with around 700B parameters and MoE layers (e.g., SparseGPT can only prune models with around 176B parameters), we use SparseGPT to prune our selected distilled models. In order to precisely analyze the collapse point of this pruning method, we run SparseGPT on each distilled model seven times to get 10%, 30%, 40%, 50%, 60%, 70%, and 80% sparsity levels.
4.3 Evaluation Metrics
Accuracy metric is used for AIME 2024, FOLIO, and Temporal Sequences. Since MuSiQue involves question answering and its answers are in a few words, we adopt exact match (EM) and F1 to measure question answering performance of different models. Without worrying inference speedup of selected compression methods, we focus on benchmarking their performance. The reason is that these methods run on different inference platforms. As a result, different optimization strategies might be adopted when we do inference, and it is hard to control the consistency of inference optimization across various platforms.
Our analysis of test-time compute involves computing the number of tokens spent on each question. We first preprocess each output by extracting reasoning tokens and sentences of final model prediction. Instead of relying on a specific tokenizer, we then treat words within each output as tokens for simplicity. We study whether there is any relationship between length and reasoning performance.
4.4 Implementation Details
Required by our quantized models, we run their inference on llama.cpp [^4] https://github.com/ggml-org/llama.cpp. While our pruned and distilled models can be deployed on various inference platforms, we use vLLM (Kwon et al., 2023) for its fast inference. In order to comprehensively analyze performance change after compression, we also evaluate R1 on our reasoning benchmarks by using DeepSeek API. Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models: maximum generation length is set to $3276832768$, temperature is set to $0.60.6$, and top-p value is set to $0.950.95$.
5 Results and Analysis
Our results and analysis aim to answer the following research questions:
  • RQ 1: How does each compression methodology compare against each other (5.1)?
  • RQ 2: Do our selected compression methodologies collapse (5.2)?
  • RQ 3: What is the recommended methodology to compress LRMs (5.1, 5.3, and 5.5)?
  • RQ 4: Is it a good sign when models spend extra amounts of test-time compute on certain questions (5.4)?
Models Accuracy
Model #Param Compression AIME 2024 FOLIO Temporal Avg MuSiQue (EM, F1)
DeepSeek-R1 671B - 73.3 76.4 99.6 83.1 (17.0, 27.51)
DeepSeek-R1 671B 2.51-bit 76.7 77.8 100.0 84.8 (17.0, 24.43)
DeepSeek-R1 671B 1.73-bit 66.7 78.3 99.6 81.5 (15.0, 22.11)
DeepSeek-R1 671B 1.58-bit 66.7 75.4 94.0 78.7 (14.0, 22.34)
R1-Distill-Llama 70B Distillation 63.3 78.8 100.0 80.7 (13.0, 21.80)
R1-Distill-Llama 70B Distillation & 50% sparse 26.7 70.9 97.2 64.9 (6.0, 12.75)
R1-Distill-Qwen 32B Distillation 66.7 82.3 100.0 83.0 (1.0, 9.38)
R1-Distill-Qwen 32B Distillation & 50% sparse 30.0 75.4 96.0 67.1 (3.0, 9.29)
Table 2: Overall results of R1 and it compressed variants. The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal. EM and F1 of MuSiQue are shown in tuples. We segment this table based on model family (e.g., R1-Distill-Llama).
5.1 Overall Performance
The overall results of R1 and its compressed variants are presented in Table 2 and 3. We only show performance of pruned R1-Distill-Llama and pruned R1-Distill-Qwen under 50% sparsity in Table 2, as it is the default sparsity level of SparseGPT and many other works (Zhang et al., 2024a; Sun et al., 2023). Table 3 displays performance of our two distilled models under various sparsity levels.
Comparing Performance of Compression Methodologies In Table 2, 2.51-bit R1 achieves the highest average accuracy scores, surpassing even the original R1. We suspect that certain security mechanisms in DeepSeek may have lowered R1’s performance, as we often receive empty responses when using its API, requiring us to prompt R1 iteratively until valid responses are obtained. Both R1-Distill-Llama and R1-Distill-Qwen have competitive average accuracy scores. According to Table 3, we see that pruning also maintains good average accuracy scores when sparsity levels are low (e.g., 30% sparsity of R1-Distill-Llama). On MuSiQue dataset, the 2.51-bit model has the same EM as R1, with a slight decrease in F1. According to scores of the closed-book setting on GPT-4o (Zhang et al., 2025), all quantized models achieve relatively strong performance on MuSiQue, with higher EM and F1 than distilled and pruned models. Therefore, this 2.51-bit model has the best overall performance than other compression methods.
Comparing Benchmark Difficulties We can make judgment about the difficulties of our accuracy-based benchmarks through Table 2. By comparing the scores of AIME 2024, FOLIO, and Temporal, we see all models struggle more on AIME 2024. This indicates that AIME 2024 is more difficult than the other two benchmarks. MuSiQue is also a difficult benchmark in terms of knowledge requirement, because its scores in Table 2 are much lower than RAG setup (Zhang et al., 2025; Jimenez Gutierrez et al., 2024). This suggests that existing LRMs lack sufficient knowledge for knowledge-intensive tasks, making RAG a more suitable approach. We demonstrate a diverse selection of reasoning benchmarks with varying difficulty levels, which is important for comprehensive evaluation.
Comparing Distilled Models based on LLaMA and Qwen When focusing on accuracy-based benchmarks, we see that R1-Distill-Qwen delivers an average 2.3% improvement over R1-Distill-Llama. This demonstrates that Qwen-32B has stronger reasoning capability than LLaMA-70B, which also aligns with the evaluation results of DeepSeek-R1 report (Guo et al., 2025). However, R1-Distill-Qwen scores significantly lower on MuSiQue, highlighting its inability to memorize detailed knowledge.
5.2 Collapse Point of Each Compression Method
We investigate whether LRMs degrade as they undergo increasing levels of compression. In Table 2, the performance of dynamically quantized models steadily declines as we move from 2.51-bit to 1.58-bit. However, we do not observe a clear collapse point, as performance declines only slightly when transitioning to a lower bit-size tier. As for distillation, comparing R1 with R1-Distill-Llama, we also do not observe a clear collapse point, although the performance drop on MuSiQue is significant. However, we see that R1-Distill-Qwen collapses on MuSiQue. With EM close to zero, it almost loses its ability on MuSiQue during distillation.
Comparing distilled models with their sparsified variants in Table 3, we find that pruned models collapse on all benchmarks at certain sparsity levels. Interestingly, their collapse point correlates to the difficulty of the benchmark. For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half. However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024. The correlation between collapse point and benchmark difficulty can also be seen on sparsified R1-Distill-Qwen. The early occurrence of collapse on AIME 2024 demonstrates that it is one of the most challenging datasets, as existing pruning works typically observe model collapse after 50% sparsity (Zhang et al., 2024a; Bai et al., 2024). Moreover, since R1-Distill-Qwen already collapses on MuSiQue, pruning it becomes meaningless for this dataset, as we observe very low scores for all sparsified R1-Distill-Qwen models on MuSiQue.
Models Accuracy
Model #Param Sparsity AIME 2024 FOLIO Temporal Avg MuSiQue (EM, F1)
R1-Distill-Llama 70B 0% 63.3 78.8 100.0 80.7 (13.0, 21.80)
R1-Distill-Llama 70B 10% 60.0 81.3 99.6 80.3 (12.0, 21.69)
R1-Distill-Llama 70B 30% 63.3 79.3 99.6 80.7 (14.0, 21.40)
R1-Distill-Llama 70B 40% 56.7 73.9 98.8 76.8 (6.0, 13.79)
R1-Distill-Llama 70B 50% 26.7 70.9 97.2 64.9 (6.0, 12.75)
R1-Distill-Llama 70B 60% 0.0 65.0 95.6 53.5 (0.0, 6.42)
R1-Distill-Llama 70B 70% 0.0 49.8 15.6 21.8 (0.0, 2.23)
R1-Distill-Llama 70B 80% 0.0 11.8 12.4 8.1 (0.0, 0.94)
R1-Distill-Qwen 32B 0% 66.7 82.3 100.0 83.0 (1.0, 9.38)
R1-Distill-Qwen 32B 10% 70.0 81.3 100.0 83.8 (5.0, 13.19)
R1-Distill-Qwen 32B 30% 56.7 81.3 100.0 79.3 (1.0, 10.47)
R1-Distill-Qwen 32B 40% 53.3 78.3 100.0 77.2 (2.0, 10.16)
R1-Distill-Qwen 32B 50% 30.0 75.4 96.0 67.1 (3.0, 9.29)
R1-Distill-Qwen 32B 60% 0.0 65.0 87.2 50.7 (0.0, 4.13)
R1-Distill-Qwen 32B 70% 0.0 32.5 19.6 17.4 (0.0, 1.72)
R1-Distill-Qwen 32B 80% 0.0 8.7 2.0 3.6 (0.0, 1.29)
Table 3: Performance of our two distilled models under various sparsity levels. The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal.
Models AIME 2024 FOLIO MuSiQue
Model Compression Short Long Ratio Short Long Ratio Short Long Ratio
DeepSeek-R1 - 88.9 33.3 5.3 83.3 63.3 8.0 (30.0, 42.8) (3.3, 10.0) 6.4
DeepSeek-R1 2.51-bit 100.0 33.3 4.9 85.0 71.7 6.5 (33.3, 41.9) (0.0, 6.9) 7.7
DeepSeek-R1 1.73-bit 88.9 22.2 4.4 86.7 73.3 4.9 (30.0, 41.6) (10.0, 23.5) 6.9
DeepSeek-R1 1.58-bit 77.8 44.4 3.8 80.0 65.0 5.0 (30.0, 41.5) (10.0, 16.6) 5.9
R1-Distill-Llama Distillation 88.9 11.1 5.4 80.0 80.0 4.5 (26.7, 38.0) (6.7, 18.4) 4.0
R1-Distill-Llama Distillation & 10% sparse 100.0 0.0 6.6 85.0 78.3 4.9 (20.0, 29.6) (6.7, 13.7) 7.4
R1-Distill-Llama Distillation & 30% sparse 88.9 11.1 5.3 85.0 76.7 5.0 (26.7, 36.9) (3.3, 12.5) 8.8
R1-Distill-Qwen Distillation 100.0 11.1 7.1 86.7 75.0 6.9 (16.7, 24.1) (16.7, 24.7) 7.1
R1-Distill-Qwen Distillation & 10% sparse 88.9 33.3 4.8 83.3 75.0 5.5 (23.3, 36.7) (3.3, 12.8) 8.1
R1-Distill-Qwen Distillation & 30% sparse 88.9 11.1 6.8 86.7 78.3 7.8 (26.7, 40.7) (3.3, 9.1) 8.8
Table 4: Analysis of test-time compute when selecting the shortest and longest 30% of responses output by each model on each benchmark. “Short” column contains performance scores of the shortest 30% of outputs from a model, while “long” column contains scores of the longest 30% of outputs. We compare the scores between “Short” and “Long” for every model and benchmark, and mark the best scores in bold. “Ratio” column represents the ratio of the average length (in number of tokens) of the longest 30% to that of the shortest 30%.
Models AIME 2024 FOLIO MuSiQue
Model Compression Short Long Ratio Short Long Ratio Short Long Ratio
DeepSeek-R1 - 100.0 16.7 7.1 82.5 57.5 10.9 (30.0, 42.4) (5.0, 12.5) 8.5
DeepSeek-R1 2.51-bit 100.0 33.3 6.6 87.5 62.5 8.6 (40.0, 50.3) (0.0, 4.5) 10.7
DeepSeek-R1 1.73-bit 100.0 33.3 5.9 87.5 67.5 6.3 (30.0, 43.7) (5.0, 15.7) 9.3
DeepSeek-R1 1.58-bit 100.0 33.3 5.6 80.0 72.5 6.2 (35.0, 46.3) (15.0, 20.0) 7.6
R1-Distill-Llama Distillation 83.3 16.7 7.0 82.5 75.0 5.7 (35.0, 45.2) (5.0, 9.2) 5.3
R1-Distill-Llama Distillation & 10% sparse 100.0 0.0 9.2 90.0 75.0 6.3 (20.0, 28.5) (5.0, 11.7) 10.7
R1-Distill-Llama Distillation & 30% sparse 100.0 16.7 7.1 90.0 75.0 6.4 (20.0, 31.0) (0.0, 7.3) 13.0
R1-Distill-Qwen Distillation 100.0 0.0 9.8 87.5 75.0 9.0 (20.0, 28.7) (20.0, 24.2) 10.4
R1-Distill-Qwen Distillation & 10% sparse 83.3 33.3 6.1 85.0 75.0 7.2 (30.0, 36.2) (0.0, 8.4) 12.0
R1-Distill-Qwen Distillation & 30% sparse 100.0 0.0 8.9 87.5 75.0 11.0 (30.0, 36.6) (5.0, 11.2) 13.4
Table 5: Analysis of test-time compute when selecting the shortest and longest 20% of responses output by each model on each benchmark. Refer to Table 4 for the meaning of each column, except that we select shortest and longest 20% instead.
5.3 Compression Impact on Knowledge and Reasoning
As discussed above, although Qwen-32B shows stronger reasoning capability than LLaMA-70B, it has significantly lower EM and F1 scores on MuSiQue. Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen-32B puts itself at a disadvantaged position. In other words, models’ parameter count affects knowledge more than reasoning. When a compression method aggressively removes the weights of an LRM , it is expected that the model’s knowledge will be more seriously affected. This phenomenon can also be seen on our quantized and pruned models. Since quantization preserves parameter count and our analysis above shows that the quantized models still retain competitive reasoning capability, it is not surprising that even the 1.58-bit model outperforms other distilled and pruned LRMs on MuSiQue. In addition, we notice that pruned R1-Distill-Llama collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024. Since MuSiQue is difficult due to its knowledge requirement, we see that pruning hurts LRMs’ knowledge memorization more than quantization.
Based on the finding that parameter count has a much greater impact on knowledge than reasoning, we recommend quantization as the LRMs compression method when some levels of models’ parametric knowledge are desired. Since both distillation and pruning involve removing weights, we recommend to choose a large student model or a low sparsity when the end task requires knowledge memorization. On the other hand, a model with strong reasoning capability does not need to be extremely large. As Qwen-32B is a stronger reasoner than LLaMA-70B, smaller models also have the potential to outperform larger ones on various reasoning tasks.
5.4 Test-time Compute
We study the behavior of R1 and its compressed variants by measuring their test-time compute. Table 4 shows the analysis of test-time compute when we select the shortest and longest 30% of responses output by each model on each benchmark. We observe that shorter model outputs consistently yield better performance across three reasoning benchmarks, with only an exception on MuSiQue. Regardless of whether an LRM is compressed, if it generates significantly more tokens for a question than other questions in the same dataset, the answer is likely to be incorrect. We exclude Temporal here, because many of the compressed R1 models achieve close to 100% accuracy. The length ratios between the longest and the shortest 30% are typically greater than 4, which indicates nontrivial length differences among model outputs.
Similarly, Table 5 shows the selection of the shortest and the longest 20% of model responses. Compared to Table 4, as we move toward the extreme, the performance gap between the shortest and longest responses becomes even larger. Both R1 and its compressed variants achieve higher scores when they spend less compute during test time. After manually checking some long responses, we notice that longer outputs tend to be more verbose and involve more backtracking in reasoning. This finding demonstrates the need to reduce verbosity to improve reasoning performance, which aligns with recent research (Zhang et al., 2024b).
5.5 Case Study
As discussed in Section 5.2, pruned models collapse on all benchmarks at certain sparsity levels. We identify a common phenomenon when a model collapses: it repeatedly generates a sentence or a chunk until reaching the maximum generation length. We show two examples of this phenomenon in Figure 2. For brevity, we omit the beginning and the end of outputs.
Figure 2: Two examples of the case when a model collapses and keeps repeating itself. These are two outputs (for a FOLIO question) from R1-Distill-Qwen at either 70% or 80% sparsity levels.
In both examples, the pruned models are repeating themselves without pushing their reasoning chains forward, which is a signal of model collapse. When R1-Distill-Qwen is pruned to 70% sparsity, we see that it can still organize a few sentences (e.g., a chunk) to repeat. But when it is pruned to 80% sparsity, it only repeats a simple sentence. This decline of linguistic capability is common when models are pruned to high sparsities. Therefore, aggressively pruning LRMs requires careful consideration.
6 Conclusion and Future Directions
In this paper, we benchmark compressed LRMs on four complex reasoning tasks: AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue. We evaluate models that are quantized, distilled, and pruned. Through MuSiQue, we find that parameter count has a much greater impact on LRMs’ knowledge than reasoning capability, which can potentially guide compression decision. Based on our study of test-time compute, we also find that shorter outputs of R1 and compressed LRMs generally achieve better performance than longer outputs. This finding highlights the need to reduce verbosity in reasoning chains.
Since compression on LRMs is relatively underexplored, future directions involve designing novel methodologies to specifically compress LRMs. The evaluation of compressed LRMs should also place greater emphasis on diverse reasoning tasks. Since DeepSeek-R1 has already been compressed through quantization and distillation, the question remains how to prune such a large model.
Appendix A Additional Details of Reasoning Benchmarks
AIME 2024555https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 (parts I and II) represents top match challenges, and its answers are integers. FOLIO [^6] https://huggingface.co/datasets/yale-nlp/FOLIO requires logical deductions to determine whether the provided conclusion is true, false, or uncertain based on premise. In Temporal Sequences [^7] https://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/temporal_sequences.json, models are asked to use a provided timeline to determine what time a person might be free to perform another activity. Since each of its questions comes with four options, we expect our models to output the index (the letter) of the selected option. The answers of MuSiQue questions are in a few words.

### v2

@@
-When Reasoning Meets Compression: Benchmarking Compressed Large Reasoning Models on Complex Reasoning Tasks
+When Reasoning Meets Compression: Understanding the Effects of LLMs Compression on Large Reasoning Models
 Abstract
-Recent open-source large reasoning models (LRMs) exhibit strong performance on complex reasoning tasks, but their large parameter count makes them prohibitively expensive for individuals.
-The compression of large language models (LLMs) offers an effective solution to reduce cost of computational resources.
-However, systematic studies on the performance of compressed LLMs in complex reasoning tasks, especially for LRMs, are lacking.
-Most works on quantization and pruning focus on preserving language modeling performance, while existing distillation works do not comprehensively benchmark student models based on reasoning difficulty or compression impact on knowledge and reasoning.
-In this paper, we benchmark compressed DeepSeek-R1 models on four different reasoning datasets (AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue), ranging from mathematical to multihop reasoning, using quantization, distillation, and pruning methods.
-We benchmark 2.51-, 1.73-, and 1.58-bit R1 models that adopt dynamic quantization.
-We also benchmark distilled R1 models that are based on LLaMA or Qwen and run SparseGPT on them to obtain various sparsity levels.
-Studying the performance and behavior of compressed LRMs, we report their performance scores and test-time compute (number of tokens spent on each question).
-Notably, using MuSiQue, we find that parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability, which can inform the choice of compression techniques.
-Through our empirical analysis of test-time compute, we find that shorter model outputs generally achieve better performance than longer ones across several benchmarks for both R1 and its compressed variants, highlighting the need for more concise reasoning chains.
+Compression methods, including quantization, distillation, and pruning, improve the computational efficiency of large reasoning models (LRMs).
+However, existing studies either fail to sufficiently compare all three compression methods on LRMs or lack in-depth interpretation analysis.
+In this paper, we investigate how the reasoning capabilities of LRMs are compromised during compression, through performance benchmarking and mechanistic interpretation.
+To uncover the effects of compression on reasoning performance, we benchmark quantized, distilled, and pruned DeepSeek-R1 models on four reasoning datasets (AIME 2024, FOLIO, Temporal Sequences, and MuSiQue).
+To precisely locate compression effects on model weights, we adapt difference of means and attribution patching techniques, focusing on the activation of every linear component in compressed LRMs, to interpret fine-grained causal relationships between weights and various reasoning capabilities.
+This fine-grained interpretation addresses a fundamental question of compression: which weights are the most important for reasoning?
+Overall, we find dynamically quantized 2.51-bit R1 reaches close-to-R1 performance.
+With empirical verification, we present three main findings that generalize across both Llama and Qwen: (1) Weight count has a greater impact on LRMs’ knowledge memorization than reasoning, highlighting the risks of pruning and distillation; (2) The MLP up projection in the final layer of distilled LRMs is one of the most important components, offering a new perspective on locating critical weights — a fundamental problem in model compression; and (3) Current quantization methods overly compress the final-layer modules and MLP gate projections, so protecting just 2% of all weights that are excessively compressed can raise average accuracy by 6.57%, greatly surpassing the state-of-the-art.
 1 Introduction
-Large reasoning models (LRMs) excel at complex reasoning tasks.
-DeepSeek-R1 (Guo et al., 2025) is the first open-source LRM to successfully replicate OpenAI’s O1-level performance.
-Trained based on DeepSeek-V3 (DeepSeek-AI et al., 2025), R1 adopts large-scale reinforcement learning without supervised fine-tuning (SFT).
-Its open-source nature facilitates research on reasoning models (Srivastava et al., 2025; Mondillo et al., 2025).
-However, due to its large size (671B total parameters), deploying it locally can be costly and even infeasible for individuals, which hinders AI democratization.
-Compression of large language models (LLMs) includes quantization, distillation, and pruning.
-Its goal is to reduce computational resources (e.g., GPU memory and disk space) and provide inference speedup.
-As the weights of vanilla LLMs are typically in 16-bit floating values (e.g.., FP16 or BF16), quantization means to convert these high-precision values to lower-precision ones.
-Representative techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization (Lin et al., 2024), and post-training quantization (Frantar et al., 2022).
-Distillation aims to transfer knowledge from a larger teacher model to a smaller student model under black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings.
-As for pruning, it aims to sparsify LLMs by introducing 0s to replace nonzero weights.
-Representative techniques include unstructured pruning (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
-Although existing quantization and pruning methods claim to preserve the performance of non-reasoning LLMs after compression such as LLaMA (Touvron et al., 2023) and OPT (Zhang et al., 2022)), exploring these methods on LRMs offers a timely research of democratizing the latest LRMs.
-Moreover, since these methods mainly adopt perplexity and simple end-task scores for evaluation, the assessment of compressed LRMs will be fundamentally different.
-Without focusing on basic linguistic capabilities (e.g., simple language understanding and generation), evaluation of compressed reasoning models needs to align with their uncompressed counterparts.
-Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider compression impact on knowledge and reasoning.
-A comprehensive evaluation that includes all popular compression methods is valuable.
-Therefore, due to the lack of compression research on LRMs, we benchmark the performance and behavior of compressed DeepSeek-R1 on various reasoning tasks in this paper.
-Our benchmarking framework is shown in Figure 1.
-We test dynamic quantization (Daniel Han & team, 2023), distillation with SFT (Guo et al., 2025), and SpraseGPT (Frantar & Alistarh, 2023) on R1 (or distilled R1) due to their popularity and effectiveness.
-Specifically, our chosen dynamic quantization reduces the MoE (Mixture of Experts) layers of R1 model to 2.51-bit, 1.73-bit, and 1.58-bit precision.
-Testing distillation with SFT, we evaluate the distilled R1 models that adopt LLaMA-70B (Meta AI, ) and Qwen-32B (Qwen et al., 2025) as students.
-Using SparseGPT, we prune the distilled R1 models to different sparsity levels.
-We select four complex reasoning benchmarks: mathematical, logical, temporal, and multihop reasoning.
-Figure 1: An overview of selected compression methods and tasks. We evaluate R1 and its compressed variants based on their performance scores and test-time compute.
-In addition to assessing model performance on selected benchmarks, we study the behavior of compressed R1 models by measuring their test-time compute (number of tokens spent on each question).
-Specifically, we compare their performance on the shortest and longest outputs.
-Our major findings are:
-  • Parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability.
-Therefore, both distillation and pruning are discouraged when the end task relies on models’ parametric knowledge.
-Otherwise, larger student models or low sparsity levels are recommended.
-  • In almost all cases, both R1 and its compressed variants achieve higher scores when they spend less compute during inference time.
-Shorter outputs tend to be more accurate than longer ones across our benchmarks.
-This highlights the need to reduce verbosity to improve reasoning performance.
-2 Related Work
-2.1 Quantization
-Quantization reduces the number of bits used to represent LLM weights, thereby lowering their precision (Srivastava et al., 2025).
-Recent survey (Zhu et al., 2024) categorizes quantization methodologies into quantization-aware training (QAT) and post-training quantization (PTQ).
-QAT requires retraining of model weights to recover performance loss during quantization while PTQ does not require retraining.
-Recent QAT includes LLM-QAT (Liu et al., 2024a) that adopts distillation to train a quantized LLM, BitDistiller (Du et al., 2024) that develops a self-distillation approach for the full-precision model to act as the teacher of its low-bit counterpart, BitNet (Wang et al., 2023) that proposes a 1-bit Transformer architecture for training LLMs from scratch, and OneBit (Xu et al., 2024) that quantizes LLM weight matrices to 1-bit from a knowledge transfer perspective.
-PTQ is more popular in terms of the number of recent publications, because there is no retraining involved.
-For example, GPTQ (Frantar et al., 2022) is a one-shot weight quantization method that uses approximate second-order information, while AWQ (Lin et al., 2024) leverages activation distribution for finding the salient weight channels to skip.
-Other PTQ methods include weight-activation quantization (Shao et al., 2024; Yao et al., 2022; Liu et al., 2023) and KV cache quantization (Hooper et al., 2024; Liu et al., 2024b).
-2.2 Distillation
-Distillation involves two settings: black-box and white-box settings.
-For black-box setting, teacher model is typically a closed-source LLM and only the outputs of teacher are available for student model.
-For white-box setting, both weights and output distribution of the teacher model are available.
-Existing black-box distillation (Huang et al., 2024; Li et al., 2024b; Ho et al., 2023; Huang et al., 2022; Li et al., 2024a) prompts the teacher model to generate a training dataset for the student to learn.
-Specifically, researchers have started to distill OpenAI’s O1 model (Huang et al., 2024), which marks the beginning of LRMs compression.
-White-box distillation allows the student model to learn from teacher’s knowledge representation.
-Works has been done to align the output distribution (Agarwal et al., 2024; Gu et al., 2024) or the hidden representation (Liang et al., 2023) between teacher and student models.
-2.3 Pruning
-There are unstructured and structured pruning.
-For unstructured pruning, individual weights are targeted, which leads to irregular sparsity structure.
-In contrast, structured pruning involves removing entire network components such as channels or layers (Zhang et al., 2024a).
-Unstructured pruning usually has better compression performance than structured pruning, while it is easier to achieve inference speedup via structured methods (Zhu et al., 2024).
-Recent unstructured pruning includes one-shot pruning (Frantar & Alistarh, 2023; Sun et al., 2023), global pruning that makes pruning decisions based on all layers (Bai et al., 2024), and domain-specific pruning (Zhang et al., 2024a).
-Structured pruning includes gradient-based  (Xia et al., 2024; Ma et al., 2023) and non-gradient-based (Ashkboos et al., 2024) methods.
-2.4 LRMs
-Trained with reinforcement learning, LRMs extends LLMs with advanced reasoning mechanisms (Besta et al., 2025).
-Popular closed-source LRMs are OpenAI’s o1-mini, o1 (OpenAI et al., 2024), and o3-mini.
-Open-source LRMs include DeepSeek-R1 and QwQ-32B-Preview (Team, 2024).
-Since quantization, white-box distillation, and pruning methods require access to model weights, they are not suitable for closed-source LRMs.
-Only black-box distillation will work on closed-source models.
-3 Background
-Scope We benchmark quantization, distillation, and pruning methodologies, because we consider them as major LLMs compression paradigms in this paper.
-Due to lower popularity, other methodologies that can potentially facilitate compression such as low-rank factorization (Sharma et al., 2024; Srebro & Jaakkola, 2003) are not elaborated.
-When selecting models for compression, we focus exclusively on LRMs.
-Recent efforts on LRMs compression As reviewed in Section 2, compression on LRMs (not LLMs) is relatively underexplored.
-Recently, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights.
-By combining four ideas (Yu et al., 2024; Dettmers et al., 2023; Ma et al., 2024; Gerganov, 2023), Unsloth quantizes DeepSeek-R1 into four variants: 2.51-, 2.22-, 1.73-, and 1.58-bit models.
-Here the number of bits (e.g., 2.51-bit) represents the precision of MoE weights, which make up 88% of the total R1 weights.
+Large reasoning models (LRMs) such as DeepSeek-R1 (Guo et al., 2025) excel at complex reasoning tasks.
+However, due to their large sizes, deploying them can be costly and even infeasible for individuals, which hinders AI democratization.
+Compression methods including quantization, distillation, and pruning reduce computational resources (e.g., GPU memory and disk space).
+Representative quantization techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization AWQ (Lin et al., 2024), and post-training quantization GPTQ (Frantar et al., 2022).
+Current distillation involves black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings.
+Representative pruning techniques include unstructured (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
+However, existing works do not sufficiently study the performance of compression method on LRMs (Liu et al., 2025; Srivastava et al., 2025; Feng et al., 2025).
+Although current quantization and pruning methods claim to preserve the performance of general-purpose LLMs, benchmarking both of them on LRMs with more reasoning-intensive datasets helps compare their collapse point.
+Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider distillation effect on knowledge and reasoning (Huang et al., 2024; Agarwal et al., 2024).
+Another research gap is the lack of interpretability of compression effects on LRMs.
+It is necessary to interpret how compression methods affect LRMs, as such analysis can reveal existing bottlenecks and provide guidance for future compression research.
+Figure 1: An overview of our pipeline.
+Left: We benchmark compressed R1 variants on various reasoning tasks.
+Right: By computing weight importance towards a specific reasoning behavior (a dot product of the steering vector and gradients with respect to an LRM’s activations), we study the compression effects on individual weight.
+We empirically verify our findings on weight importance by selectively quantizing or protecting a module to test its importance.
+Therefore, due to the lack of compression works on LRMs, we study this fundamental research question: How are the reasoning capabilities of LRMs compressed during compression?
+We answer it from two perspectives — performance benchmarking and mechanistic interpretation.
+We first benchmark compressed DeepSeek-R1 on various reasoning tasks to investigate how model compression affects performance.
+We test dynamic quantization (Daniel Han & team, 2023), distillation with supervised fine-tuning (SFT) (Guo et al., 2025), SparseGPT (Frantar & Alistarh, 2023), AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) on R1 (or distilled R1).
+Then, we apply mechanistic interpretability to quantify weight contribution towards four core reasoning capabilities of LRMs: backtracking, uncertainty estimation, example testing, and adding knowledge.
+By focusing on the activation of every linear component in compressed LRMs, we adapt difference of means (Arditi et al., 2024) to extract steering vectors and attribution patching (Syed et al., 2023) to compute weight importance.
+Unlike previous analysis (Venhoff et al., 2025) that only measures layer-wise weight contribution, our weight importance scores offer more fine-grained interpretation of weight contribution, addressing the fundamental compression question of locating important weights.
+By comparing weight importance scores between compressed LRMs and original LRMs, we quantify the effects of distillation and quantization on model weights [^1] Our interpretation code is at https://github.com/psunlpgroup/Compression-Effects..
+Our analysis framework is shown in Figure 1.
+With empirical verification, our key findings are summarized below:
+  • Weight count has a greater impact on LRMs’ knowledge memorization than their reasoning capabilities, highlighting the compression effects of pruning and distillation.
+Thus, both distillation and pruning are discouraged when tasks require LRMs’ parametric knowledge.
+  • The mlp.up_proj in the final layer of R1 distilled models emerges as one of the most important model components, addressing a core concern in pruning and quantization literature: identifying critical weights.
+Quantizing only this matrix to 3-bit reduces the average accuracy by 16.3%.
+  • Final-layer modules, along with the mlp.gate_proj of R1 distilled Llama and Qwen, are overly compressed by popular quantization methods, highlighting the need for greater attention to preserving their weight precision.
+A successful protection of only final-layer MLP modules could raise average accuracy by 6.57%, with gains of up to 23.17% over the state-of-the-art quantization.
+2 Problem Formulation
+2.1 Background
+As discussed in Section 1, compression on LRMs (not LLMs) is relatively underexplored. We conduct a thorough literature review in Appendix B.
+Bottlenecks on evaluation.
+Few quantization or pruning methodologies have sufficiently demonstrated effectiveness on LRMs.
+Current works evaluate quantization and pruning performance primarily using perplexity and simple end tasks, such as the EleutherAI evaluation harness (Gao et al., 2024) and commonsense reasoning.
+However, quantized or pruned LRMs should be assessed on more complex reasoning tasks with varying difficulty levels.
+For distillation, although recent works tend to test on more challenging reasoning tasks (compared to other compression literature) such as GSM8K (Cobbe et al., 2021), it is unclear how the compression of LRMs affects models’ parametric knowledge and reasoning capability.
+Some of them do not comprehensively select diverse reasoning benchmarks (Agarwal et al., 2024).
+Our benchmarking aims to address these bottlenecks.
+Bottlenecks on in-depth analysis.
+The lack of interpretability of compression effects on LRMs is a key bottleneck of in-depth analysis for compressed LRMs.
+Being able to interpret the difference between original and compressed LRMs offers a new way to analyze the effects of compression.
+As a result, better compression approaches can be developed.
+A recent work (Venhoff et al., 2025) interprets several R1 distilled LRMs, but their focus is not on understanding compression effects.
+Recent efforts.
+Recent benchmarking (Liu et al., 2025) and survey (Feng et al., 2025; Srivastava et al., 2025) papers have begun to evaluate compressed LRMs on more complex reasoning datasets, but they all lack in-depth interpretation of compression effects and do not comprehensively compare different compression strategies.
+As for compressed LRMs, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights.
 DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation.
-Current bottlenecks Very few quantization or pruning works have demonstrated effectiveness on LRMs, as LRMs have only recently emerged.
-Moreover, evaluation of quantized or pruned LRMs will be different from quantized or pruned LLMs.
-Current works evaluate quantization and pruning performance primarily using perplexity and simple end tasks, such as the EleutherAI evaluation harness (Gao et al., 2024) and commonsense reasoning.
-However, compressed LRMs should be assessed on more complex reasoning tasks with varying difficulty levels.
-As for distillation, although recent distillation works tend to test on more challenging reasoning tasks (compared to other compression literature) such as GSM8K (Cobbe et al., 2021), it is unclear how the compression of LRMs affects models’ parametric knowledge and reasoning capability.
-Some of them do not comprehensively select diverse reasoning benchmarks.
-Our benchmarking framework aims to address these bottlenecks.
-#Questions #Pass Answer Type Metric Knowledge required?
-AIME 2024 30 1 Integer Accuracy False
-FOLIO 203 1 True/False/Uncertain Accuracy False
-Temporal 250 1 (A)/(B)/(C)/(D) Accuracy False
-MuSiQue 100 1 A few words (EM, F1) True
-Table 1: Statistics of selected benchmarks: AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue. We use exact match (EM) and F1 to measure model performance on MuSiQue.
-4 Experiment Setup
-4.1 Reasoning Benchmarks
-To address the evaluation bottleneck discussed in Section 3, we select four different reasoning benchmarks with varying levels of difficulty: AIME 2024 (Mathematical Association of America, ) for mathematical reasoning, FOLIO (Han et al., 2024) for logical reasoning, Temporal Sequences of BIG-Bench Hard (Suzgun et al., 2022) for temporal reasoning, and MuSiQue (Trivedi et al., 2022) for multihop reasoning.
-Table 1 summarizes their statistics.
-For each model, we do a single pass on every benchmark.
-Unlike other benchmarks where we select all questions, we randomly sample 100 out of 1000 from MuSiQue, as it is much larger.
-This pool of 1000 questions is the same as SiReRAG (Zhang et al., 2025) and HippoRAG (Jimenez Gutierrez et al., 2024).
-Since MuSiQue requires knowledge memorization besides multihop reasoning, it is popular under retrieval-augmented generation (RAG) setup.
-Instead of RAG, we follow a closed-book setting (directly prompting LRMs to get final answers) to test models’ parametric knowledge.
-In this case, unlike other benchmarks that test only reasoning capability, we use MuSiQue to evaluate both reasoning and knowledge memorization.
-Additional details of selected benchmarks are specified in Appendix A.
-4.2 Selected Compression Methods
-We decide to compress DeepSeek-R1 due to it large parameter count.
-For quantization, we select 2.51-, 1.73-, and 1.58-bit models [^1] https://huggingface.co/unsloth/DeepSeek-R1-GGUF by Unsloth.
-These dynamically quantized models are among the most widely used quantization methods on R1.
-As for distillation, we select the two largest distilled models that accompany R1: DeepSeek-R1-Distill-Llama-70B [^2] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-32B [^3] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-32B.
-Since we do not find a pruning method designed for models with around 700B parameters and MoE layers (e.g., SparseGPT can only prune models with around 176B parameters), we use SparseGPT to prune our selected distilled models.
-In order to precisely analyze the collapse point of this pruning method, we run SparseGPT on each distilled model seven times to get 10%, 30%, 40%, 50%, 60%, 70%, and 80% sparsity levels.
-4.3 Evaluation Metrics
+Our interpretation analysis aims to demystify the effects of LLMs compression on LRMs, providing a systematic understanding of existing compressed LRMs.
+2.2 Mechanistic Interpretation
+For our interpretation analysis, we target four core reasoning behaviors following an existing work (Venhoff et al., 2025): backtracking, uncertainty estimation, example testing, and adding knowledge.
+We prompt GPT-4o to locate token sequences of each behavior from the output tokens of our LRMs.
+To interpret different compression strategies, we adapt difference of means and attribution patching by computing the activation of every linear module in each layer.
+This allows us to compute the causal relationship between each weight matrix and our target reasoning behaviors.
+Difference of Means.
+To compute the numerical representation in activation space of each reasoning behavior, we adapt difference of means method (Venhoff et al., 2025; Arditi et al., 2024) to extract the steering vector $\mathbf{u}_{m\ell}^{c}$ for each linear module $m$ at layer $\ell$ for behavior $c$:
+$\mathbf{u}_{m\ell}^{c}=\frac{1}{\lvert\mathcal{D}_{+}\rvert}\sum_{s_{i}^{c}\in\mathcal{D}_{+}}\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})-\frac{1}{\lvert\mathcal{D}_{-}\rvert}\sum_{s_{j}\in\mathcal{D}_{-}}\overline{\mathbf{a}}_{m\ell}(s_{j}),\quad\text{with}\quad\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})=\frac{1}{\lvert s_{i}^{c}\rvert}\sum_{t\in s_{i}^{c}}\mathbf{a}_{m\ell}(t)$
+where $s_{i}^{c}$ denotes the token sequence corresponding to a specific reasoning behavior $c$ along with its five preceding tokens as output by an LRM, $s_{j}$ is the token sequence of the entire LRM output (prompt and output tokens), $\mathcal{D}_{+}$ is the set of output instances containing at least one token sequence labeled with $c$, $\mathcal{D}_{-}$ is the set of all output instances, $\mathbf{a}_{m\ell}(t)$ is the activation of module $m$ at layer $\ell$ at token $t$, $\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})$ is the average of $\mathbf{a}_{m\ell}(t)$ across all tokens in $s_{i}^{c}$, and similarly, $\overline{\mathbf{a}}_{m\ell}(s_{j})$ is the average of $\mathbf{a}_{m\ell}(t)$ across all tokens in $s_{j}$.
+We then normalize $\mathbf{u}_{m\ell}^{c}$ to $\tilde{\mathbf{u}}_{m\ell}^{c}$: $\tilde{\mathbf{u}}_{m\ell}^{c}=\mathbf{u}_{m\ell}^{c}\cdot\frac{\left\lVert\overline{\mathbf{a}}_{m\ell}^{\mathrm{all}}\right\rVert_{2}}{\left\lVert\mathbf{u}_{m\ell}^{c}\right\rVert_{2}}$ where $\overline{\mathbf{a}}_{m\ell}^{\mathrm{all}}$ denotes the mean activation across all tokens in $\mathcal{D}_{-}$.
+Attribution Patching.
+To find the causally relevant LRMs components with respect to each reasoning behavior, we adapt attribution patching (Syed et al., 2023) method to compute the importance score $\mathbf{I}_{m\ell}^{c}$ of each linear module.
+$\mathbf{I}_{m\ell}^{c}\approx\frac{1}{\lvert\mathcal{D}_{+}\rvert}\left|\sum_{s_{i}^{c}\in\mathcal{D}_{+}}\big(\tilde{\mathbf{u}}_{m\ell}^{\,c}\big)^{\top}\,\frac{\partial}{\partial\mathbf{a}_{m\ell}}\,\mathcal{L}\!\left(s_{i}^{c}\right)\right|$
+where $\mathcal{L}\!\left(s_{i}^{c}\right)$ is the cross-entropy loss of $s_{i}^{c}$.
+A higher $\mathbf{I}_{m\ell}^{c}$ means a stronger causal relationship between $c$ and the linear module $m$ at layer $\ell$, helping us locate the most important weights responsible for reasoning capabilities (a fundamental problem for quantization and pruning works).
+2.3 Decoding Compression Effects
+To decode compression effects, we compute the relative importance $\mathbf{RI}_{m\ell}^{c}$ of each weight matrix ($\mathbf{I}_{m\ell}^{c}$ divided by $\sum_{m}\sum_{\ell}\mathbf{I}_{m\ell}^{c}$) and track how it changes because of compression (importance shift).
+Specifically, we measure the change of $\mathbf{RI}_{m\ell}^{c}$ from R1 distilled Llama-8B to original meta-llama/Llama-3.1-8B to understand distillation effect (Section 4).
+Likewise, the importance shift from the R1 distilled models to their quantized versions indicates quantization effect (Section 5).
+For the R1 distilled models, we also compute the $\mathbf{I}_{m\ell}^{c}$ of their weights to complement our findings on the distillation effect.
+We hypothesize that the importance shift should be minimal in the ideal case, as a compressed LRM should remain as close as possible to its original counterpart (the more reasoning-capable model).
+When visualizing the importance shift from an LRM to its compressed variant (or from a distilled model to its backbone), we only consider decreases in $\mathbf{RI}_{m\ell}^{c}$.
+By definition, the relative importance of each weight matrix is normalized to sum to one, so any increase in relative importance necessarily compensates for decreases elsewhere.
+Since it is more informative to track cases where the $\mathbf{RI}_{m\ell}^{c}$ of a more reasoning-capable model decreases (e.g., when the reasoning capability of a weight matrix is diminished), we set all increases in relative importance to zero.
+2.4 Scope
+We study three major LLMs compression paradigms, distillation, quantization, and pruning, making our scope comprehensive enough for investigating the effects of diverse compression methods.
+For distillation, we select four R1 distilled models: DeepSeek-R1-Distill Llama-70B, Qwen-32B, Llama-8B, and Qwen-7B.
+For quantization, we select 2.51-, 1.73-, and 1.58-bit models by Unsloth [^2] https://huggingface.co/unsloth/DeepSeek-R1-GGUF as the choices of quantized R1 due to their popularity.
+We also evaluate AWQ Lin et al.
+(2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) as reproducible state-of-the-art quantization methods designed for relatively smaller LLMs (e.g., the R1 distilled models).
+Specifically, we use all four methods to perform 4-bit quantization, and use GPTQ, GPTAQ, and ANY3 for 3-bit quantization as well, since many AWQ implementations do not support 3-bit.
+For pruning, we run SparseGPT Frantar & Alistarh (2023) on our two largest distilled models.
+We run interpretaion analysis on linear modules of all layers within LRMs.
+2.5 Evaluation Setup
+We select four reasoning datasets with varying levels of difficulty: AIME 2024 (Mathematical Association of America, ) for mathematical reasoning, FOLIO (Han et al., 2024) for logical reasoning, Temporal Sequences of BIG-Bench Hard (Suzgun et al., 2022) for temporal reasoning, and MuSiQue (Trivedi et al., 2022) for multihop reasoning.
+Since MuSiQue requires knowledge memorization besides multihop reasoning, we follow a closed-book setting (directly prompting LRMs to get final answers) to evaluate both reasoning and knowledge retention capabilities.
+Additional details of benchmarks, along with Table 5 that shows their statistics, are specified in Appendix C.
 Accuracy metric is used for AIME 2024, FOLIO, and Temporal Sequences.
-Since MuSiQue involves question answering and its answers are in a few words, we adopt exact match (EM) and F1 to measure question answering performance of different models.
-Without worrying inference speedup of selected compression methods, we focus on benchmarking their performance.
-The reason is that these methods run on different inference platforms.
-As a result, different optimization strategies might be adopted when we do inference, and it is hard to control the consistency of inference optimization across various platforms.
-Our analysis of test-time compute involves computing the number of tokens spent on each question.
-We first preprocess each output by extracting reasoning tokens and sentences of final model prediction.
-Instead of relying on a specific tokenizer, we then treat words within each output as tokens for simplicity.
-We study whether there is any relationship between length and reasoning performance.
-4.4 Implementation Details
-Required by our quantized models, we run their inference on llama.cpp [^4] https://github.com/ggml-org/llama.cpp.
-While our pruned and distilled models can be deployed on various inference platforms, we use vLLM (Kwon et al., 2023) for its fast inference.
-In order to comprehensively analyze performance change after compression, we also evaluate R1 on our reasoning benchmarks by using DeepSeek API.
-Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models: maximum generation length is set to $3276832768$, temperature is set to $0.60.6$, and top-p value is set to $0.950.95$.
-5 Results and Analysis
-Our results and analysis aim to answer the following research questions:
-  • RQ 1: How does each compression methodology compare against each other (5.1)?
-  • RQ 2: Do our selected compression methodologies collapse (5.2)?
-  • RQ 3: What is the recommended methodology to compress LRMs (5.1, 5.3, and 5.5)?
-  • RQ 4: Is it a good sign when models spend extra amounts of test-time compute on certain questions (5.4)?
+We adopt exact match (EM) and F1 for MuSiQue.
+For each model (except R1 and those dynamically quantized LRMs), we run it three times and report its average scores to mitigate performance variability.
+Implementation details are in Appendix D.
+3 Compression Effects on Reasoning Performance
+3.1 Overall Performance
+The overall performance of R1 and its compressed variants are in Table 1.
+We show the performance of pruned R1-Distill-Llama-70B and R1-Distill-Qwen-32B under 50% sparsity in Table 1, as it is the default sparsity level of current works (Zhang et al., 2024a; Sun et al., 2023).
+Additional analysis of test-time compute and model collapse behavior are in Appendices G and H.
+Table 1: Benchmark performance of R1 and its compressed variants.
+All four benchmark scores are averaged over three passes, except the rows marked with .
+Avg denotes the average scores shown in AIME 2024, FOLIO, and Temporal columns.
+We segment this table based on model families and mark the highest scores within each model family in bold.
 Models Accuracy
@@
 DeepSeek-R1 671B 1.58-bit 66.7 75.4 94.0 78.7 (14.0, 22.34)
-R1-Distill-Llama 70B Distillation 63.3 78.8 100.0 80.7 (13.0, 21.80)
-R1-Distill-Llama 70B Distillation & 50% sparse 26.7 70.9 97.2 64.9 (6.0, 12.75)
-R1-Distill-Qwen 32B Distillation 66.7 82.3 100.0 83.0 (1.0, 9.38)
-R1-Distill-Qwen 32B Distillation & 50% sparse 30.0 75.4 96.0 67.1 (3.0, 9.29)
-Table 2: Overall results of R1 and it compressed variants.
-The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal.
-EM and F1 of MuSiQue are shown in tuples.
-We segment this table based on model family (e.g., R1-Distill-Llama).
-5.1 Overall Performance
-The overall results of R1 and its compressed variants are presented in Table 2 and 3.
-We only show performance of pruned R1-Distill-Llama and pruned R1-Distill-Qwen under 50% sparsity in Table 2, as it is the default sparsity level of SparseGPT and many other works (Zhang et al., 2024a; Sun et al., 2023).
-Table 3 displays performance of our two distilled models under various sparsity levels.
-Comparing Performance of Compression Methodologies In Table 2, 2.51-bit R1 achieves the highest average accuracy scores, surpassing even the original R1.
-We suspect that certain security mechanisms in DeepSeek may have lowered R1’s performance, as we often receive empty responses when using its API, requiring us to prompt R1 iteratively until valid responses are obtained.
-Both R1-Distill-Llama and R1-Distill-Qwen have competitive average accuracy scores.
-According to Table 3, we see that pruning also maintains good average accuracy scores when sparsity levels are low (e.g., 30% sparsity of R1-Distill-Llama).
-On MuSiQue dataset, the 2.51-bit model has the same EM as R1, with a slight decrease in F1.
-According to scores of the closed-book setting on GPT-4o (Zhang et al., 2025), all quantized models achieve relatively strong performance on MuSiQue, with higher EM and F1 than distilled and pruned models.
-Therefore, this 2.51-bit model has the best overall performance than other compression methods.
-Comparing Benchmark Difficulties We can make judgment about the difficulties of our accuracy-based benchmarks through Table 2.
-By comparing the scores of AIME 2024, FOLIO, and Temporal, we see all models struggle more on AIME 2024.
-This indicates that AIME 2024 is more difficult than the other two benchmarks.
-MuSiQue is also a difficult benchmark in terms of knowledge requirement, because its scores in Table 2 are much lower than RAG setup (Zhang et al., 2025; Jimenez Gutierrez et al., 2024).
+R1-Distill-Llama 70B Distillation 65.6 79.8 99.9 81.8 (13.3, 21.57)
+R1-Distill-Llama 70B Distillation & 50% sparse 23.3 71.6 97.6 64.2 (6.7, 13.49)
+R1-Distill-Llama 70B Distillation & 4-bit AWQ 63.4 78.5 99.3 80.4 (10.7, 19.23)
+R1-Distill-Llama 70B Distillation & 4-bit GPTQ 66.7 77.0 99.9 81.2 (10.3, 18.17)
+R1-Distill-Llama 70B Distillation & 4-bit GPTAQ 64.4 78.8 99.6 80.9 (12.0, 21.57)
+R1-Distill-Llama 70B Distillation & 3-bit GPTQ 46.7 71.8 99.3 72.6 (4.7, 11.92)
+R1-Distill-Llama 70B Distillation & 3-bit GPTAQ 54.4 77.3 99.7 77.1 (5.7, 13.21)
+R1-Distill-Qwen 32B Distillation 64.4 82.3 99.9 82.2 (2.7, 10.95)
+R1-Distill-Qwen 32B Distillation & 50% sparse 25.6 75.1 97.9 66.2 (2.3, 9.01)
+R1-Distill-Qwen 32B Distillation & 4-bit AWQ 67.8 82.3 99.1 83.1 (3.3, 10.28)
+R1-Distill-Qwen 32B Distillation & 4-bit GPTQ 68.9 80.6 99.6 83.0 (4.0, 11.78)
+R1-Distill-Qwen 32B Distillation & 4-bit GPTAQ 63.3 81.5 99.7 81.5 (2.7, 11.88)
+R1-Distill-Qwen 32B Distillation & 4-bit ANY4 68.9 78.0 99.7 82.2 (5.7, 12.68)
+R1-Distill-Qwen 32B Distillation & 3-bit GPTQ 44.4 74.2 98.9 72.5 (4.0, 11.55)
+R1-Distill-Qwen 32B Distillation & 3-bit GPTAQ 45.6 77.5 99.5 74.2 (2.3, 9.18)
+R1-Distill-Qwen 32B Distillation & 3-bit ANY3 53.3 82.6 99.9 78.6 (3.7, 10.27)
+R1-Distill-Llama 8B Distillation 42.2 71.9 81.5 65.2 (0.0, 4.43)
+R1-Distill-Llama 8B Distillation & 4-bit AWQ 47.8 68.0 84.0 66.6 (0.3, 5.05)
+R1-Distill-Llama 8B Distillation & 4-bit GPTQ 42.2 66.2 65.9 58.1 (0.3, 4.68)
+R1-Distill-Llama 8B Distillation & 4-bit GPTAQ 40.0 66.4 69.3 58.6 (0.0, 3.73)
+R1-Distill-Llama 8B Distillation & 4-bit ANY4 41.1 68.5 88.7 66.1 (0.0, 3.54)
+R1-Distill-Llama 8B Distillation & 3-bit GPTQ 11.1 65.0 67.3 47.8 (0.0, 2.89)
+R1-Distill-Llama 8B Distillation & 3-bit GPTAQ 7.8 65.5 57.2 43.5 (0.0, 3.45)
+R1-Distill-Llama 8B Distillation & 3-bit ANY3 3.3 50.1 34.9 29.4 (0.7, 2.35)
+R1-Distill-Qwen 7B Distillation 46.7 78.0 75.6 66.8 (0.0, 3.57)
+R1-Distill-Qwen 7B Distillation & 4-bit AWQ 46.6 75.5 74.9 65.7 (0.0, 3.14)
+R1-Distill-Qwen 7B Distillation & 4-bit GPTQ 38.9 72.9 70.3 60.7 (1.0, 4.27)
+R1-Distill-Qwen 7B Distillation & 4-bit GPTAQ 47.8 74.4 67.7 63.3 (0.0, 3.96)
+R1-Distill-Qwen 7B Distillation & 4-bit ANY4 47.8 75.6 77.1 66.8 (0.0, 3.05)
+R1-Distill-Qwen 7B Distillation & 3-bit GPTQ 17.8 65.7 31.7 38.4 (0.0, 3.12)
+R1-Distill-Qwen 7B Distillation & 3-bit GPTAQ 24.4 64.5 48.7 45.9 (0.0, 3.06)
+R1-Distill-Qwen 7B Distillation & 3-bit ANY3 32.2 69.3 30.1 43.9 (0.0, 3.89)
+Comparing Compression Strategies.
+In Table 1, the 2.51-bit R1 achieves the highest average accuracy overall, since it has the smallest compression ratio.
+Both R1 distilled Llama-70B and Qwen-32B reach close-to-R1 accuracy scores.
+On MuSiQue, the 2.51-bit R1 also achieves performance close to original R1.
+Therefore, 2.51-bit R1 has the best overall performance than other compression strategies.
+Although R1 may be over-parameterized, a compression method with a smaller ratio can still offer advantages over methods with higher compression ratios.
+In contrast, pruning only 50% of the weights causes significant degradation, rendering the pruned LRMs unusable.
+Thus, we do not interpret the effect of pruning in the later sections.
+As for all distillation-only models, Qwen delivers stronger reasoning performance than Llama (Appendix E).
+Comparing Benchmark Difficulties.
+Comparing the scores using R1 distilled Llama-70B on AIME 2024, FOLIO, and Temporal, we see the largest score decrease on AIME 2024.
+This indicates that AIME 2024 is more challenging than the other two accuracy-based benchmarks.
+MuSiQue is also difficult in terms of knowledge requirement, because its scores in Table 1 are much lower than RAG (retrieval-augmented generation) setup (Zhang et al., 2025).
 This suggests that existing LRMs lack sufficient knowledge for knowledge-intensive tasks, making RAG a more suitable approach.
-We demonstrate a diverse selection of reasoning benchmarks with varying difficulty levels, which is important for comprehensive evaluation.
-Comparing Distilled Models based on LLaMA and Qwen When focusing on accuracy-based benchmarks, we see that R1-Distill-Qwen delivers an average 2.3% improvement over R1-Distill-Llama.
-This demonstrates that Qwen-32B has stronger reasoning capability than LLaMA-70B, which also aligns with the evaluation results of DeepSeek-R1 report (Guo et al., 2025).
-However, R1-Distill-Qwen scores significantly lower on MuSiQue, highlighting its inability to memorize detailed knowledge.
-5.2 Collapse Point of Each Compression Method
+Takeaway 3.1 for Overall Performance Considering over-parameterization, methods with smaller compression ratios can still offer advantages over those with higher compression ratios.
+Regardless of whether compression is applied, LRMs lack sufficient knowledge for knowledge-intensive tasks.
+3.2 Collapse Point
 We investigate whether LRMs degrade as they undergo increasing levels of compression.
-In Table 2, the performance of dynamically quantized models steadily declines as we move from 2.51-bit to 1.58-bit.
-However, we do not observe a clear collapse point, as performance declines only slightly when transitioning to a lower bit-size tier.
-As for distillation, comparing R1 with R1-Distill-Llama, we also do not observe a clear collapse point, although the performance drop on MuSiQue is significant.
-However, we see that R1-Distill-Qwen collapses on MuSiQue.
-With EM close to zero, it almost loses its ability on MuSiQue during distillation.
-Comparing distilled models with their sparsified variants in Table 3, we find that pruned models collapse on all benchmarks at certain sparsity levels.
-Interestingly, their collapse point correlates to the difficulty of the benchmark.
-For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half.
-However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024.
-The correlation between collapse point and benchmark difficulty can also be seen on sparsified R1-Distill-Qwen.
-The early occurrence of collapse on AIME 2024 demonstrates that it is one of the most challenging datasets, as existing pruning works typically observe model collapse after 50% sparsity (Zhang et al., 2024a; Bai et al., 2024).
-Moreover, since R1-Distill-Qwen already collapses on MuSiQue, pruning it becomes meaningless for this dataset, as we observe very low scores for all sparsified R1-Distill-Qwen models on MuSiQue.
+In Table 1, the performance of dynamically quantized LRMs steadily declines as we move from 2.51 to 1.58-bit, but we do not observe a clear collapse point.
+All 4-bit AWQ, GPTQ, GPTAQ, and ANY4 reach performance similar to their unquantized counterparts, which shows the effectiveness of existing 4-bit quantization on LRMs.
+However, 3-bit GPTQ, GPTAQ, and ANY3 display signs of collapse, indicating bottlenecks of current 3-bit quantization.
+GPTAQ and ANY4 are newer than AWQ and GPTQ, and they achieve similar performance on 4-bit LRMs.
+Based on average accuracy, it is noteworthy that GPTAQ surpasses GPTQ on 3-bit LRMs for three out of four distilled models.
+Regarding distillation, R1 distilled Llama-8B and Qwen-7B achieve the lowest accuracy among all distillation-only models.
+Only R1 distilled Llama-70B yields decent MuSiQue scores.
+Table 2: Performance of two distilled models under various sparsity levels. We report the one-pass scores for all models in this table.
 Models Accuracy
@@
 R1-Distill-Qwen 32B 80% 0.0 8.7 2.0 3.6 (0.0, 1.29)
-Table 3: Performance of our two distilled models under various sparsity levels. The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal.
+Table 2 displays performance of our two distilled models under various sparsity levels.
+Comparing distilled models with their sparsified variants, we find the precise collapse points of our pruned LRMs.
+Interestingly, their collapse points correlate to the benchmark difficulty.
+For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half.
+However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024.
+The correlation between collapse point and benchmark difficulty can also be seen on the sparsified Qwen.
+Takeaway 3.2 for Collapse Point Collapse point correlates with benchmark difficulty.
+On hard benchmarks, 3-bit quantization and pruning with 50% sparsity or higher still have substantial room for improvement.
+3.3 Compression Impact on Knowledge and Reasoning
+In Table 2, although Qwen demonstrates stronger reasoning capabilities than Llama, it has significantly lower EM and F1 scores on MuSiQue.
+Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen puts itself at a disadvantaged position.
+In other words, models’ parameter count affects knowledge more than reasoning.
+In addition, we notice pruned R1-Distill-Llama-70B collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024.
+This shows pruning hurts LRMs’ knowledge memorization more than quantization.
+When a compression method aggressively removes the weights of an LRM, it is expected that the model’s knowledge will be more seriously affected.
+This phenomenon can also be seen on our dynamically quantized models in Table 1.
+Since quantization preserves parameter count and our analysis above shows that many quantized models still retain competitive reasoning capability, quantization is recommended on knowledge-intensive tasks.
+Takeaway 3.3 for Compression Impact on Knowledge and Reasoning Pruning and distillation compress knowledge retention more than reasoning capabilities.
+4 Distillation Effect on Weights
+To study the effect of distillation on weights, we compute $\mathbf{I}_{m\ell}^{c}$ of two distilled R1 models and further measure the change of $\mathbf{RI}_{m\ell}^{c}$ as discussed in Section 2.3.
+4.1 Locating Important Weights
+The left part of Figure 2 presents the weight importance of R1 distilled Llama-8B in four heatmaps, each corresponding to a reasoning behavior.
+We observe that the final layer houses several most important linear modules across all four behaviors, with the highest value located at up_proj.
+Therefore, the up_proj in the final layer (32_up) stands out as the most important component.
+Interestingly, this finding generalizes to R1 distilled Qwen-7B, as we also observe this up_proj outlier in the final layer of Qwen in Figure 4.
+Notably, our finding complements a recent analysis (Shao & Wu, 2025), which claims the most important module for reasoning is o_proj.
+Since identifying important weights is a core research problem of compression methodologies, our finding is valuable for future works.
+Takeaway 4.1 for Locating Important Weights Distillation makes up_proj in the final layer as the most important module for reasoning behaviors, as observed in both R1 distilled Llama and Qwen models.
+Figure 2: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Llama-8B (left) and change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Llama-3.1-8B (right).
+Each heatmap displays scores of importance (or importance shift) of every module at each layer, providing a fine-grained analysis of weight contributions to the corresponding reasoning capability.
+On the right, increases in $\mathbf{RI}_{m\ell}^{c}$ are set to 0, as they only offset decreases elsewhere as discussed in 2.3.
+Every cluster of 4 side-by-side heatmaps (including those displayed below) follow the same scaling to show the precise magnitude of each weight module.
+Figure 3: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant.
+Table 3: Accuracy after selectively quantizing a single component of R1 distilled Llama-8B (e.g., 1_up means to only quantize the up_proj in the first layer) to 3-bit.
+Ranking of a component is based its $\sum_{c}\mathbf{I}_{m\ell}$, so “2nd col” refers to second place within its column across all four heatmaps (each column consists of 7 linear modules of a layer).
+“1st overall” means the global highest ranking.
+Quantized Component Rank AIME 2024 FOLIO Temporal Avg
+32_up 1st overall 20.0 63.1 63.6 48.9
+32_gate 2nd col 33.3 62.1 67.2 54.2
+32_v last col 43.3 68.0 79.6 63.6
+31_up 2nd row 33.3 70.0 64.4 55.9
+1_up last row 6.7 64.5 80.4 50.5
+4.2 Validating Importance Scores
+We validate Section 4.1 by applying 3-bit round-to-nearest quantization to either 32_up or a component sharing its column or row in the heatmaps, then measuring the resulting accuracy drop (Table 3).
+The more important a component is, the greater the accuracy drop when it is quantized.
+Specifically, we select four additional component candidates: the second- and last-ranked modules among the seven in the final layer, and the second- and last-ranked layers across all 32 layers of the up-projection.
+We see that 32_up yields the lowest average accuracy, which clearly demonstrates the validity of our findings in Section 4.1.
+It is quite salient that quantizing only this matrix (merely 0.7% of all weights) reduces the average accuracy by 16.3%.
+The component rank generally correlates with the accuracy drop, except for 1_up, which incurs the lowest accuracy on AIME 2024.
+4.3 Importance Shift via Distillation
+Since R1-Distill-Llama-8B is fine-tuned based on Llama-3.1-8B, we compute the change of $\mathbf{RI}_{m\ell}^{c}$ to visualize distillation effect in the right part of Figure 2 (for Llama) and 5 (for Qwen).
+Both parts of Figure 2 exhibit similar patterns (e.g., most outliers are in the final layer), indicating that the important weights of the distilled model are primarily the result of distillation with SFT, while the original Llama’s weight values play little role in shaping its reasoning capabilities.
+Thus, distillation effect is quite powerful in transforming a non-reasoning LLM into an LRM.
+For Qwen, Figures 4 and 5 also show similar patterns, so the utility of distillation generalizes to Qwen as well.
+Takeaway 4.3 for Importance Shift via Distillation Important weights of the R1 distilled models are mainly the result of the distillation effect.
+5 Quantization Effect on Weights
+For quantization effect on weights, we analyze the decrease of importance shift during quantization.
+5.1 Locating Quantization Effect
+We show heatmaps to visualize the importance shift from R1 distilled Llama-8B to its 4-bit AWQ quantized variant in Figure 3.
+Across all four heatmaps, we observe a reduction in the significance of the gate projections in the middle layers (e.g., layer 9 to 23), suggesting that AWQ may overly compress these modules.
+Moreover, most linear modules in the final layer are compressed to the greatest extent, which shows the drawback of AWQ.
+Since 32_up is the most important module as discussed in Section 4.2 and its importance shift is little for uncertainty estimation and example testing capabilities, AWQ successfully preserves its significance on these two behaviors.
+However, for backtracking and adding knowledge capabilities, AWQ is not effective at maintaining its importance.
+In Figure 6, we visualize the importance shift from R1 distilled Qwen-7B to its 4-bit AWQ quantized version.
+We also see a shift in the importance of the gate projections on Qwen, but this shift mainly occurs in the early layers (e.g., layer 1 to 10).
+On Qwen, AWQ does not preserve the importance of 32_up across all four reasoning capabilities, and it also overly compresses 32_k on two capabilities.
+As another popular method, we interpret the effect of 4-bit GPTQ in Figure 7.
+On R1 distilled Llama-8B, we observe similar quantization effect as AWQ, since GPTQ also overly compresses final-layer modules and the gate projections in the middle layers.
+Their commonality demonstrates the generality of the bottlenecks we identified in existing quantization methods.
+Takeaway 5 for Quantization Effect on Weights State-of-the-art quantization methods fail to preserve the importance of the MLP gate projections and the final layer, which is a key bottleneck of performance improvement.
+5.2 Validating Quantization Effect
+Table 4: Performance of 3-bit AWQ and selectively protecting the MLP modules in the final layer.
+Model Compression Full-Precision Anywhere? AIME 2024 FOLIO Temporal Avg MuSiQue
+R1-Distill-Llama-8B 3-bit AWQ - 10.0 59.6 68.4 46.0 (0.0, 3.50)
+R1-Distill-Llama-8B 3-bit AWQ Final-layer MLP 16.7 67.0 74.0 52.57 (1.0, 3.62)
+To validate our findings about the bottleneck of current quantization, we design a simple protection mechanism using a mixed precision fashion.
+We run two versions of 3-bit AWQ in Table 4.
+In the first version, we run AWQ with their default calibration data.
+Since we know AWQ overly compresses the MLP modules in the final layer, we then choose to protect them by changing their quantized weights to their original values in 16-bit.
+If they are truly important yet not well protected by AWQ, our protection mechanism should offer a significant improvement.
+Based on the discussion in Sections 3.1 and 3.2, we perform 3-bit quantization, since it will be the focus of future works.
+We see our selective protection boosts 3-bit AWQ on all benchmarks, with an average accuracy improvement of 6.57%.
+This is particularly significant given that only about 2% of all weights remain in 16-bit.
+This mixed precision model outperforms all 3-bit quantization baselines in Table 1 by at least 4.77% in average accuracy, with gains of up to 23.17%.
+Therefore, our findings are demonstrated with an indication of substantial room for further improvement.
+Note that our protection provides relatively marginal increase on MuSiQue, as the weight count stays the same (Section 3.3).
+6 Conclusion and Future Directions
+We study the effects of LLMs compression on LRMs and present key findings for further imrpving LRMs compression.
+Future compression works are encouraged to consider the protection of MLP up projection in the final layer.
+The excessive compression of current quantization methods on MLP gate projections and final-layer modules highlights the need for preserving these weight modules.
+Appendix A Use of LLMs
+To improve the overall clarity of our writing, we used ChatGPT-4o and ChatGPT-5 via OpenAI’s web interface to polish a small fraction of sentences.
+LLMs were not used in any steps of the research ideation process.
+To ensure correctness and precision, we carefully reviewed and adapted all LLM-generated content before incorporating it into our writing.
+Appendix B Related Work
+Our literature review is conducted over existing compression methodologies (quantization, distillation, and pruning) and recent LRMs.
+B.1 Quantization
+Quantization reduces the number of bits used to represent LLM weights, thereby lowering their precision (Srivastava et al., 2025).
+Recent survey (Zhu et al., 2024) categorizes quantization methodologies into quantization-aware training (QAT) and post-training quantization (PTQ).
+QAT requires retraining of model weights to recover performance loss during quantization while PTQ does not require retraining.
+Recent QAT includes LLM-QAT (Liu et al., 2024a) that adopts distillation to train a quantized LLM, BitDistiller (Du et al., 2024) that develops a self-distillation approach for the full-precision model to act as the teacher of its low-bit counterpart, BitNet (Wang et al., 2023) that proposes a 1-bit Transformer architecture for training LLMs from scratch, and OneBit (Xu et al., 2024) that quantizes LLM weight matrices to 1-bit from a knowledge transfer perspective.
+PTQ is more popular in terms of the number of recent publications, because there is no retraining involved.
+For example, GPTQ (Frantar et al., 2022) and GPTAQ (Li et al., 2025) are one-shot weight quantization methods that use approximate second-order information, while AWQ (Lin et al., 2024) leverages activation distribution for finding the salient weight channels to skip.
+Other PTQ methods include weight-activation quantization (Shao et al., 2024; Yao et al., 2022; Liu et al., 2023) and KV cache quantization (Hooper et al., 2024; Liu et al., 2024b).
+B.2 Distillation
+Distillation involves two settings: black-box and white-box settings.
+For black-box setting, the teacher model is typically a closed-source LLM and only the outputs of teacher are available for the student model.
+For white-box setting, both weights and output distribution of the teacher model are available.
+Existing black-box distillation (Huang et al., 2024; Li et al., 2024b; Ho et al., 2023; Huang et al., 2022; Li et al., 2024a) prompts the teacher model to generate a training dataset for the student to learn.
+Specifically, researchers have started to distill OpenAI’s O1 model (Huang et al., 2024), which marks the beginning of LRMs compression.
+White-box distillation allows the student model to learn from teacher’s knowledge representation.
+Works has been done to align the output distribution (Agarwal et al., 2024; Gu et al., 2024) or the hidden representation (Liang et al., 2023) between teacher and student models.
+B.3 Pruning
+There are unstructured and structured pruning.
+For unstructured pruning, individual weights are targeted, which leads to irregular sparsity structure.
+In contrast, structured pruning involves removing entire network components such as channels or layers (Zhang et al., 2024a).
+Unstructured pruning usually has better compression performance than structured pruning, while it is easier to achieve inference speedup via structured methods (Zhu et al., 2024).
+Recent unstructured pruning includes one-shot pruning (Frantar & Alistarh, 2023; Sun et al., 2023), global pruning that makes pruning decisions based on all layers (Bai et al., 2024), and domain-specific pruning (Zhang et al., 2024a).
+Structured pruning includes gradient-based  (Xia et al., 2024; Ma et al., 2023) and non-gradient-based (Ashkboos et al., 2024) methods.
+B.4 LRMs
+Trained with reinforcement learning, LRMs extends LLMs with advanced reasoning mechanisms (Besta et al., 2025).
+Popular closed-source LRMs are OpenAI’s o1-mini, o1 (OpenAI et al., 2024), and o3-mini.
+Open-source LRMs include DeepSeek-R1 and QwQ-32B-Preview (Team, 2024).
+Since quantization, white-box distillation, and pruning methods require access to model weights, they are not suitable for closed-source LRMs.
+Only black-box distillation will work on closed-source models.
+Table 5: Dataset statistics of selected reasoning benchmarks.
+Size Answer Type Metric Knowledge
+AIME 2024 30 Integer Accuracy False
+FOLIO 203 True/False/Uncertain Accuracy False
+Temporal 250 (A)/(B)/(C)/(D) Accuracy False
+MuSiQue 100 A few words (EM, F1) True
+Appendix C Additional Details of Reasoning Benchmarks
+Table 5 shows the statistics of our selected benchmarks.
+AIME 2024333https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 (parts I and II) represents top match challenges, and its answers are integers.
+FOLIO [^4] https://huggingface.co/datasets/yale-nlp/FOLIO requires logical deductions to determine whether the provided conclusion is true, false, or uncertain based on premise.
+In Temporal Sequences [^5] https://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/temporal_sequences.json, models are asked to use a provided timeline to determine what time a person might be free to perform another activity.
+Since each of its questions comes with four options, we expect our models to output the index (the letter) of the selected option.
+Since MuSiQue involves question answering and its answers are in a few words, we adopt exact match (EM) and F1.
+We randomly sample 100 questions out of 1000 from MuSiQue for our benchmarking analysis.
+Appendix D Implementation Details
+We run the dynamically quantized models on llama.cpp [^6] https://github.com/ggml-org/llama.cpp based on their requirement.
+We run all other distilled, pruned, and quantized models on vLLM (Kwon et al., 2023) for its fast inference.
+In order to comprehensively analyze performance change after compression, we also evaluate R1 on our reasoning benchmarks by using DeepSeek API.
+Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models during inference: maximum generation length is set to $32768$, temperature is set to $0.6$, and top-p value is set to $0.95$.
+We use AutoAWQ [^7] https://github.com/casper-hansen/AutoAWQ as the AWQ implementation for inference due to its speed advantage (vLLM support), while the original AWQ code [^8] https://github.com/mit-han-lab/llm-awq is used to generate the pseudo-quantized R1 distilled Llama-8B for our analysis in Section 5.
+We focus on analyzing the effect of compression methods on performance and thus do not consider inference speedup.
+The reason is that these methods run on different inference platforms, so it is hard to control the consistency of inference optimization across various platforms.
+Appendix E Comparing Distilled Models
+On accuracy-based benchmarks of Table 1, we see that R1 distilled Qwen-32B delivers an average 0.4% improvement over Llama-70 and R1 distilled Qwen-7B delivers an average 1.6% improvement over Llama-8B.
+Although these two Qwen models have less weights, Qwen delivers stronger reasoning performance than Llama.
+This phenomenon aligns with DeepSeek report (Guo et al., 2025).
+However, R1-Distill-Qwen-32B scores significantly lower than R1-Distill-Llama-70B on MuSiQue, highlighting its worse ability of memorization.
+Appendix F Additional Visualization of Weight Importance and Importance Shift
+All additional figures in Appendix are thoroughly discussed in the main content.
+Figure 4 shows the weight importance of DeepSeek-R1-Distill-Qwen-7B across four heatmaps, each corresponding to a specific target reasoning behavior.
+Figure 5 displays the change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Qwen2.5-Math-7B.
+To decode the quantization effect on Qwen, Figure 6 shows the change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Qwen-7B to its 4-bit AWQ variant.
+Similarly, Figure 7 shows the change of $\mathbf{RI}_{m\ell}^{c}$ from R1 distilled Llama-8B to its 4-bit GPTQ quantized variant.
+Figure 4: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Qwen-7B.
+Figure 5: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Qwen2.5-Math-7B (the backbone model).
+Figure 6: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Qwen-7B to its 4-bit AWQ variant.
+Figure 7: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit GPTQ quantized variant.
+Appendix G Test-time Compute
+We study the behavior of R1 and some of its compressed variants by measuring their test-time compute.
+Table 6 shows the analysis of test-time compute when we select the shortest and longest 30% of responses output by each model on each benchmark.
+We observe that shorter model outputs consistently yield better performance across three reasoning benchmarks, with only an exception on MuSiQue.
+Regardless of whether an LRM is compressed, if it generates significantly more tokens for a question than other questions in the same dataset, the answer is likely to be incorrect.
+We exclude Temporal here, because many of the compressed R1 models achieve close to 100% accuracy.
+The length ratios between the longest and the shortest 30% are typically greater than 4, which indicates nontrivial length differences among model outputs.
+Similarly, Table 7 shows the selection of the shortest and the longest 20% of model responses.
+Compared to Table 6, as we move toward the extreme, the performance gap between the shortest and longest responses becomes even larger.
+Both R1 and its compressed variants achieve higher scores when they spend less compute during test time.
+After manually checking long responses, we notice that longer outputs tend to be more verbose and involve more backtracking in reasoning.
+This finding demonstrates the need to reduce verbosity to improve reasoning performance, which aligns with recent research (Zhang et al., 2024b).
+Table 6: Analysis of test-time compute when selecting the shortest and longest 30% of responses output by each model on each benchmark.
+“Short” column contains performance scores of the shortest 30% of outputs from a model, while “long” column contains scores of the longest 30% of outputs.
+We compare the scores between “Short” and “Long” for every model and benchmark, and mark the best scores in bold.
+“Ratio” column represents the ratio of the average length (in number of tokens) of the longest 30% to that of the shortest 30%.
 Models AIME 2024 FOLIO MuSiQue
@@
 R1-Distill-Qwen Distillation & 30% sparse 88.9 11.1 6.8 86.7 78.3 7.8 (26.7, 40.7) (3.3, 9.1) 8.8
-Table 4: Analysis of test-time compute when selecting the shortest and longest 30% of responses output by each model on each benchmark.
-“Short” column contains performance scores of the shortest 30% of outputs from a model, while “long” column contains scores of the longest 30% of outputs.
-We compare the scores between “Short” and “Long” for every model and benchmark, and mark the best scores in bold.
-“Ratio” column represents the ratio of the average length (in number of tokens) of the longest 30% to that of the shortest 30%.
+Table 7: Analysis of test-time compute when selecting the shortest and longest 20% of responses output by each model on each benchmark.
+Refer to Table 6 for the meaning of each column, except that we select shortest and longest 20% instead.
 Models AIME 2024 FOLIO MuSiQue
@@
 R1-Distill-Qwen Distillation & 30% sparse 100.0 0.0 8.9 87.5 75.0 11.0 (30.0, 36.6) (5.0, 11.2) 13.4
-Table 5: Analysis of test-time compute when selecting the shortest and longest 20% of responses output by each model on each benchmark.
-Refer to Table 4 for the meaning of each column, except that we select shortest and longest 20% instead.
-5.3 Compression Impact on Knowledge and Reasoning
-As discussed above, although Qwen-32B shows stronger reasoning capability than LLaMA-70B, it has significantly lower EM and F1 scores on MuSiQue.
-Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen-32B puts itself at a disadvantaged position.
-In other words, models’ parameter count affects knowledge more than reasoning.
-When a compression method aggressively removes the weights of an LRM , it is expected that the model’s knowledge will be more seriously affected.
-This phenomenon can also be seen on our quantized and pruned models.
-Since quantization preserves parameter count and our analysis above shows that the quantized models still retain competitive reasoning capability, it is not surprising that even the 1.58-bit model outperforms other distilled and pruned LRMs on MuSiQue.
-In addition, we notice that pruned R1-Distill-Llama collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024.
-Since MuSiQue is difficult due to its knowledge requirement, we see that pruning hurts LRMs’ knowledge memorization more than quantization.
-Based on the finding that parameter count has a much greater impact on knowledge than reasoning, we recommend quantization as the LRMs compression method when some levels of models’ parametric knowledge are desired.
-Since both distillation and pruning involve removing weights, we recommend to choose a large student model or a low sparsity when the end task requires knowledge memorization.
-On the other hand, a model with strong reasoning capability does not need to be extremely large.
-As Qwen-32B is a stronger reasoner than LLaMA-70B, smaller models also have the potential to outperform larger ones on various reasoning tasks.
-5.4 Test-time Compute
-We study the behavior of R1 and its compressed variants by measuring their test-time compute.
-Table 4 shows the analysis of test-time compute when we select the shortest and longest 30% of responses output by each model on each benchmark.
-We observe that shorter model outputs consistently yield better performance across three reasoning benchmarks, with only an exception on MuSiQue.
-Regardless of whether an LRM is compressed, if it generates significantly more tokens for a question than other questions in the same dataset, the answer is likely to be incorrect.
-We exclude Temporal here, because many of the compressed R1 models achieve close to 100% accuracy.
-The length ratios between the longest and the shortest 30% are typically greater than 4, which indicates nontrivial length differences among model outputs.
-Similarly, Table 5 shows the selection of the shortest and the longest 20% of model responses.
-Compared to Table 4, as we move toward the extreme, the performance gap between the shortest and longest responses becomes even larger.
-Both R1 and its compressed variants achieve higher scores when they spend less compute during test time.
-After manually checking some long responses, we notice that longer outputs tend to be more verbose and involve more backtracking in reasoning.
-This finding demonstrates the need to reduce verbosity to improve reasoning performance, which aligns with recent research (Zhang et al., 2024b).
-5.5 Case Study
+Appendix H Case Study
 As discussed in Section 3.2, pruned models collapse on all benchmarks at certain sparsity levels.
@@
 For brevity, we omit the beginning and the end of outputs.
-Figure 2: Two examples of the case when a model collapses and keeps repeating itself. These are two outputs (for a FOLIO question) from R1-Distill-Qwen at either 70% or 80% sparsity levels.
+Figure 8: Two examples of the case when a model collapses and keeps repeating itself.
+These are two outputs (for a FOLIO question) from DeepSeek-R1-Distill-Qwen-32B at either 70% or 80% sparsity levels.
 In both examples, the pruned models are repeating themselves without pushing their reasoning chains forward, which is a signal of model collapse.
-When R1-Distill-Qwen is pruned to 70% sparsity, we see that it can still organize a few sentences (e.g., a chunk) to repeat.
+When R1-Distill-Qwen-32B is pruned to 70% sparsity, we see that it can still organize a few sentences (e.g., a chunk) to repeat.
 But when it is pruned to 80% sparsity, it only repeats a simple sentence.
@@
 Therefore, aggressively pruning LRMs requires careful consideration.
-6 Conclusion and Future Directions
-In this paper, we benchmark compressed LRMs on four complex reasoning tasks: AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue.
-We evaluate models that are quantized, distilled, and pruned.
-Through MuSiQue, we find that parameter count has a much greater impact on LRMs’ knowledge than reasoning capability, which can potentially guide compression decision.
-Based on our study of test-time compute, we also find that shorter outputs of R1 and compressed LRMs generally achieve better performance than longer outputs.
-This finding highlights the need to reduce verbosity in reasoning chains.
-Since compression on LRMs is relatively underexplored, future directions involve designing novel methodologies to specifically compress LRMs.
-The evaluation of compressed LRMs should also place greater emphasis on diverse reasoning tasks.
-Since DeepSeek-R1 has already been compressed through quantization and distillation, the question remains how to prune such a large model.
-Appendix A Additional Details of Reasoning Benchmarks
-AIME 2024555https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 (parts I and II) represents top match challenges, and its answers are integers.
-FOLIO [^6] https://huggingface.co/datasets/yale-nlp/FOLIO requires logical deductions to determine whether the provided conclusion is true, false, or uncertain based on premise.
-In Temporal Sequences [^7] https://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/temporal_sequences.json, models are asked to use a provided timeline to determine what time a person might be free to perform another activity.
-Since each of its questions comes with four options, we expect our models to output the index (the letter) of the selected option.
-The answers of MuSiQue questions are in a few words.

### v3

@@
-When Reasoning Meets Compression: Understanding the Effects of LLMs Compression on Large Reasoning Models
+arXiv:2504.02010v3[cs.LG] 02 Mar 2026
 Abstract
@@
 Overall, we find dynamically quantized 2.51-bit R1 reaches close-to-R1 performance.
-With empirical verification, we present three main findings that generalize across both Llama and Qwen: (1) Weight count has a greater impact on LRMs’ knowledge memorization than reasoning, highlighting the risks of pruning and distillation; (2) The MLP up projection in the final layer of distilled LRMs is one of the most important components, offering a new perspective on locating critical weights — a fundamental problem in model compression; and (3) Current quantization methods overly compress the final-layer modules and MLP gate projections, so protecting just 2% of all weights that are excessively compressed can raise average accuracy by 6.57%, greatly surpassing the state-of-the-art.
+With empirical verification, we present three main findings that generalize across both R1 and non-R1 LRMs: (1) Weight count has a greater impact on LRMs’ knowledge memorization than reasoning, highlighting the risks of pruning and distillation; (2) The MLP up projection in the final layer of distilled LRMs is one of the most important components, offering a new perspective on locating critical weights - a fundamental problem in model compression; and (3) Current quantization methods overly compress the final-layer modules and MLP gate projections, so protecting just 2% of all weights that are excessively compressed can raise average accuracy by 6.57%, greatly surpassing the state-of-the-art.
 1 Introduction
@@
 Compression methods including quantization, distillation, and pruning reduce computational resources (e.g., GPU memory and disk space).
-Representative quantization techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization AWQ (Lin et al., 2024), and post-training quantization GPTQ (Frantar et al., 2022).
+Representative quantization techniques include dynamic quantization by Unsloth (Daniel Han and team, 2023), activation-aware quantization AWQ (Lin et al., 2024), and post-training quantization GPTQ (Frantar et al., 2022).
 Current distillation involves black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings.
-Representative pruning techniques include unstructured (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
-However, existing works do not sufficiently study the performance of compression method on LRMs (Liu et al., 2025; Srivastava et al., 2025; Feng et al., 2025).
+Representative pruning techniques include unstructured (Zhang et al., 2024a; Frantar and Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
+However, existing works do not sufficiently study the performance of compression method on LRMs (Liu et al., 2025a; Srivastava et al., 2025; Feng et al., 2025).
 Although current quantization and pruning methods claim to preserve the performance of general-purpose LLMs, benchmarking both of them on LRMs with more reasoning-intensive datasets helps compare their collapse point.
@@
 Figure 1: An overview of our pipeline.
-Left: We benchmark compressed R1 variants on various reasoning tasks.
-Right: By computing weight importance towards a specific reasoning behavior (a dot product of the steering vector and gradients with respect to an LRM’s activations), we study the compression effects on individual weight.
+On the left, we benchmark compressed R1 variants on various reasoning tasks.
+On the right, by computing weight importance towards a specific reasoning behavior (a dot product of the steering vector and gradients with respect to an LRM’s activations), we study the compression effects on individual weight matrices.
 We empirically verify our findings on weight importance by selectively quantizing or protecting a module to test its importance.
 Therefore, due to the lack of compression works on LRMs, we study this fundamental research question: How are the reasoning capabilities of LRMs compressed during compression?
-We answer it from two perspectives — performance benchmarking and mechanistic interpretation.
-We first benchmark compressed DeepSeek-R1 on various reasoning tasks to investigate how model compression affects performance.
-We test dynamic quantization (Daniel Han & team, 2023), distillation with supervised fine-tuning (SFT) (Guo et al., 2025), SparseGPT (Frantar & Alistarh, 2023), AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) on R1 (or distilled R1).
+We answer it from two perspectives: performance benchmarking and mechanistic interpretation.
+In the main text, we first benchmark compressed DeepSeek-R1 on various reasoning tasks to investigate how model compression affects performance.
+We test dynamic quantization (Daniel Han and team, 2023), distillation with supervised fine-tuning (SFT) (Guo et al., 2025), SparseGPT (Frantar and Alistarh, 2023), AlphaPruning (Lu et al., 2024), AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi and Johnson, 2025) on R1 (or distilled R1).
 Then, we apply mechanistic interpretability to quantify weight contribution towards four core reasoning capabilities of LRMs: backtracking, uncertainty estimation, example testing, and adding knowledge.
@@
 Unlike previous analysis (Venhoff et al., 2025) that only measures layer-wise weight contribution, our weight importance scores offer more fine-grained interpretation of weight contribution, addressing the fundamental compression question of locating important weights.
-By comparing weight importance scores between compressed LRMs and original LRMs, we quantify the effects of distillation and quantization on model weights [^1] Our interpretation code is at https://github.com/psunlpgroup/Compression-Effects..
-Our analysis framework is shown in Figure 1.
-With empirical verification, our key findings are summarized below:
+By comparing weight importance between compressed and original LRMs, we quantify the effects of distillation, quantization, and pruning on model weights [^1] Our interpretation code is at https://github.com/psunlpgroup/Compression-Effects..
+Our analysis framework is in Figure 1.
+With empirical verification, our key findings are summarized below for better understanding and improving LRMs compression (these findings also generalize to non-R1 families):
   • Weight count has a greater impact on LRMs’ knowledge memorization than their reasoning capabilities, highlighting the compression effects of pruning and distillation.
@@
   • The mlp.up_proj in the final layer of R1 distilled models emerges as one of the most important model components, addressing a core concern in pruning and quantization literature: identifying critical weights.
-Quantizing only this matrix to 3-bit reduces the average accuracy by 16.3%.
+Quantizing only this matrix to 3-bit reduces the average accuracy by 16.3%, which validates its high importance.
   • Final-layer modules, along with the mlp.gate_proj of R1 distilled Llama and Qwen, are overly compressed by popular quantization methods, highlighting the need for greater attention to preserving their weight precision.
 A successful protection of only final-layer MLP modules could raise average accuracy by 6.57%, with gains of up to 23.17% over the state-of-the-art quantization.
+This key finding also applies to current pruning methods.
 2 Problem Formulation
@@
 Recent efforts.
-Recent benchmarking (Liu et al., 2025) and survey (Feng et al., 2025; Srivastava et al., 2025) papers have begun to evaluate compressed LRMs on more complex reasoning datasets, but they all lack in-depth interpretation of compression effects and do not comprehensively compare different compression strategies.
-As for compressed LRMs, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights.
+Recent benchmarking (Liu et al., 2025a) and survey (Feng et al., 2025; Srivastava et al., 2025) papers have begun to evaluate compressed LRMs on more complex reasoning datasets, but they all lack in-depth interpretation of compression effects and do not comprehensively compare different compression strategies.
+As for compressed LRMs, Unsloth (Daniel Han and team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights.
 DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation.
@@
 We prompt GPT-4o to locate token sequences of each behavior from the output tokens of our LRMs.
+Our annotation dataset consists of 120 instances drawn from the four benchmark datasets (30 instances from each), so our interpretability analysis spans four different task types and difficulties.
+Annotation robustness of GPT-4o is demonstrated in Appendix G.
 To interpret different compression strategies, we adapt difference of means and attribution patching by computing the activation of every linear module in each layer.
@@
 Since it is more informative to track cases where the $\mathbf{RI}_{m\ell}^{c}$ of a more reasoning-capable model decreases (e.g., when the reasoning capability of a weight matrix is diminished), we set all increases in relative importance to zero.
+Additional justification of only visualizing the decreases is in Appendix H.
 2.4 Scope
@@
 We also evaluate AWQ Lin et al.
-(2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) as reproducible state-of-the-art quantization methods designed for relatively smaller LLMs (e.g., the R1 distilled models).
+(2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi and Johnson, 2025) as reproducible state-of-the-art quantization methods designed for relatively smaller LLMs (e.g., the R1 distilled models).
 Specifically, we use all four methods to perform 4-bit quantization, and use GPTQ, GPTAQ, and ANY3 for 3-bit quantization as well, since many AWQ implementations do not support 3-bit.
-For pruning, we run SparseGPT Frantar & Alistarh (2023) on our two largest distilled models.
-We run interpretaion analysis on linear modules of all layers within LRMs.
+For pruning, we run SparseGPT Frantar and Alistarh (2023) and AlphaPruning (Lu et al., 2024) on several distilled models.
+Inside AlphaPruning, SparseGPT is applied to prune the layers after layerwise compression ratios are computed.
+We run interpretation analysis on linear modules of all layers within LRMs.
 2.5 Evaluation Setup
@@
 3 Compression Effects on Reasoning Performance
+In this section, we examine the effects of compression on reasoning performance.
+Interestingly, our key findings from the following sections also generalize to non-R1 model families, as elaborated in Appendix J.
 3.1 Overall Performance
@@
 We show the performance of pruned R1-Distill-Llama-70B and R1-Distill-Qwen-32B under 50% sparsity in Table 1, as it is the default sparsity level of current works (Zhang et al., 2024a; Sun et al., 2023).
-Additional analysis of test-time compute and model collapse behavior are in Appendices G and H.
+Additional analysis of test-time compute and model collapse behavior are in Appendices O and P.
 Table 1: Benchmark performance of R1 and its compressed variants.
@@
 R1-Distill-Llama 70B Distillation 65.6 79.8 99.9 81.8 (13.3, 21.57)
-R1-Distill-Llama 70B Distillation & 50% sparse 23.3 71.6 97.6 64.2 (6.7, 13.49)
+R1-Distill-Llama 70B Distillation & 50% SparseGPT 23.3 71.6 97.6 64.2 (6.7, 13.49)
+R1-Distill-Llama 70B Distillation & 50% AlphaPruning 26.7 74.2 97.7 66.2 (5.3, 12.39)
 R1-Distill-Llama 70B Distillation & 4-bit AWQ 63.4 78.5 99.3 80.4 (10.7, 19.23)
@@
 R1-Distill-Qwen 32B Distillation 64.4 82.3 99.9 82.2 (2.7, 10.95)
-R1-Distill-Qwen 32B Distillation & 50% sparse 25.6 75.1 97.9 66.2 (2.3, 9.01)
+R1-Distill-Qwen 32B Distillation & 50% SparseGPT 25.6 75.1 97.9 66.2 (2.3, 9.01)
 R1-Distill-Qwen 32B Distillation & 4-bit AWQ 67.8 82.3 99.1 83.1 (3.3, 10.28)
@@
 R1-Distill-Llama 8B Distillation 42.2 71.9 81.5 65.2 (0.0, 4.43)
+R1-Distill-Llama 8B Distillation & 30% AlphaPruning 41.1 68.9 82.1 64.1 (0.3, 4.51)
+R1-Distill-Llama 8B Distillation & 50% AlphaPruning 6.7 61.7 79.6 49.3 (0.0, 2.95)
 R1-Distill-Llama 8B Distillation & 4-bit AWQ 47.8 68.0 84.0 66.6 (0.3, 5.05)
@@
 In contrast, pruning only 50% of the weights causes significant degradation, rendering the pruned LRMs unusable.
-Thus, we do not interpret the effect of pruning in the later sections.
+Thus, we choose to interpret the effect of pruning with greater caution and specify the details in Appendix I.
 As for all distillation-only models, Qwen delivers stronger reasoning performance than Llama (Appendix E).
@@
 Only R1 distilled Llama-70B yields decent MuSiQue scores.
-Table 2: Performance of two distilled models under various sparsity levels. We report the one-pass scores for all models in this table.
+Table 2 displays performance of our two distilled models under various sparsity levels.
+Comparing distilled models with their sparsified variants, we find the precise collapse points of our pruned LRMs.
+Interestingly, their collapse points correlate to the benchmark difficulty.
+For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half.
+However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024.
+The correlation between collapse point and benchmark difficulty can also be seen on the sparsified Qwen.
+Takeaway 3.2 for Collapse Point Collapse point correlates with benchmark difficulty.
+On hard benchmarks, 3-bit quantization and pruning with 50% sparsity or higher still have substantial room for improvement.
+Table 2: Performance of two distilled models under various sparsity levels of SparseGPT. We report the one-pass scores for all models in this table.
 Models Accuracy
@@
 R1-Distill-Qwen 32B 80% 0.0 8.7 2.0 3.6 (0.0, 1.29)
-Table 2 displays performance of our two distilled models under various sparsity levels.
-Comparing distilled models with their sparsified variants, we find the precise collapse points of our pruned LRMs.
-Interestingly, their collapse points correlate to the benchmark difficulty.
-For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half.
-However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024.
-The correlation between collapse point and benchmark difficulty can also be seen on the sparsified Qwen.
-Takeaway 3.2 for Collapse Point Collapse point correlates with benchmark difficulty.
-On hard benchmarks, 3-bit quantization and pruning with 50% sparsity or higher still have substantial room for improvement.
 3.3 Compression Impact on Knowledge and Reasoning
@@
 In other words, models’ parameter count affects knowledge more than reasoning.
-In addition, we notice pruned R1-Distill-Llama-70B collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024.
-This shows pruning hurts LRMs’ knowledge memorization more than quantization.
+In addition, we notice that pruned R1-Distill-Llama-70B collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024.
+This shows that pruning hurts LRMs’ knowledge memorization more than quantization.
 When a compression method aggressively removes the weights of an LRM, it is expected that the model’s knowledge will be more seriously affected.
@@
 Since quantization preserves parameter count and our analysis above shows that many quantized models still retain competitive reasoning capability, quantization is recommended on knowledge-intensive tasks.
+Additional experiments on knowledge retrieval are specified in Appendix L.
 Takeaway 3.3 for Compression Impact on Knowledge and Reasoning Pruning and distillation compress knowledge retention more than reasoning capabilities.
@@
 4.1 Locating Important Weights
-The left part of Figure 2 presents the weight importance of R1 distilled Llama-8B in four heatmaps, each corresponding to a reasoning behavior.
+The upper part of Figure 2 presents the weight importance of R1 distilled Llama-8B in four heatmaps, each corresponding to a reasoning behavior.
 We observe that the final layer houses several most important linear modules across all four behaviors, with the highest value located at up_proj.
@@
 Interestingly, this finding generalizes to R1 distilled Qwen-7B, as we also observe this up_proj outlier in the final layer of Qwen in Figure 4.
-Notably, our finding complements a recent analysis (Shao & Wu, 2025), which claims the most important module for reasoning is o_proj.
+Notably, our finding complements a recent analysis (Shao and Wu, 2025), which claims the most important module for reasoning is o_proj.
 Since identifying important weights is a core research problem of compression methodologies, our finding is valuable for future works.
+Additional diagnosis based on AWQ also reinforces our finding on the last-layer up projection (Appendix M).
 Takeaway 4.1 for Locating Important Weights Distillation makes up_proj in the final layer as the most important module for reasoning behaviors, as observed in both R1 distilled Llama and Qwen models.
-Figure 2: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Llama-8B (left) and change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Llama-3.1-8B (right).
+Figure 2: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Llama-8B (upper) and change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Llama-3.1-8B (lower).
 Each heatmap displays scores of importance (or importance shift) of every module at each layer, providing a fine-grained analysis of weight contributions to the corresponding reasoning capability.
-On the right, increases in $\mathbf{RI}_{m\ell}^{c}$ are set to 0, as they only offset decreases elsewhere as discussed in 2.3.
+For the lower part, increases in $\mathbf{RI}_{m\ell}^{c}$ are set to 0, as they only offset decreases elsewhere as discussed in 2.3.
 Every cluster of 4 side-by-side heatmaps (including those displayed below) follow the same scaling to show the precise magnitude of each weight module.
 Figure 3: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant.
+Justification of only showing the decrease of $\mathbf{RI}_{m\ell}^{c}$ in the main content is specified in Appendix H.
 Table 3: Accuracy after selectively quantizing a single component of R1 distilled Llama-8B (e.g., 1_up means to only quantize the up_proj in the first layer) to 3-bit.
@@
 “1st overall” means the global highest ranking.
+We provide additional validation on AIME 2024 scores in Appendix N.
 Quantized Component Rank AIME 2024 FOLIO Temporal Avg
@@
 4.3 Importance Shift via Distillation
-Since R1-Distill-Llama-8B is fine-tuned based on Llama-3.1-8B, we compute the change of $\mathbf{RI}_{m\ell}^{c}$ to visualize distillation effect in the right part of Figure 2 (for Llama) and 5 (for Qwen).
-Both parts of Figure 2 exhibit similar patterns (e.g., most outliers are in the final layer), indicating that the important weights of the distilled model are primarily the result of distillation with SFT, while the original Llama’s weight values play little role in shaping its reasoning capabilities.
+Since R1-Distill-Llama-8B is fine-tuned based on Llama-3.1-8B, we compute the change of $\mathbf{RI}_{m\ell}^{c}$ to visualize distillation effect in the lower part of Figure 2 (for Llama) and 5 (for Qwen).
+Both parts of Figure 2 exhibit similar patterns (e.g., most outliers are in the final layer), indicating that the important weights of the distilled model are primarily the result of distillation with SFT, while the original Llama’s weight values play little role in shaping its reasoning capabilities (we elaborate on $\mathbf{I}_{m\ell}^{c}$ of Llama-3.1-8B in Appendix K).
 Thus, distillation effect is quite powerful in transforming a non-reasoning LLM into an LRM.
 For Qwen, Figures 4 and 5 also show similar patterns, so the utility of distillation generalizes to Qwen as well.
+In conclusion, distillation effect explains why the final-layer up projection is one of the most important LRMs components for reasoning.
 Takeaway 4.3 for Importance Shift via Distillation Important weights of the R1 distilled models are mainly the result of the distillation effect.
@@
 For quantization effect on weights, we analyze the decrease of importance shift during quantization.
+Pruning effect based on AlphaPruning appears very similar to quantization effect and is specified in Appendix I.
 5.1 Locating Quantization Effect
@@
 6 Conclusion and Future Directions
-We study the effects of LLMs compression on LRMs and present key findings for further imrpving LRMs compression.
+We study the effects of LLMs compression on LRMs and present key findings for further improving LRMs compression.
+These findings generalize across both R1 and non-R1 models.
 Future compression works are encouraged to consider the protection of MLP up projection in the final layer.
-The excessive compression of current quantization methods on MLP gate projections and final-layer modules highlights the need for preserving these weight modules.
+The excessive compression of current quantization and pruning methods on MLP gate projections and final-layer modules highlights the need for better preserving these weight modules.
+Acknowledgments
+This work was supported by NSF IIS-2338418. We thank Yanchi Liu for his valuable feedback on the project.
 Appendix A Use of LLMs
@@
 Appendix B Related Work
-Our literature review is conducted over existing compression methodologies (quantization, distillation, and pruning) and recent LRMs.
+Our literature review is conducted over existing compression methodologies (quantization, distillation, and pruning), LRMs, and recent module-wise compression.
 B.1 Quantization
@@
 Unstructured pruning usually has better compression performance than structured pruning, while it is easier to achieve inference speedup via structured methods (Zhu et al., 2024).
-Recent unstructured pruning includes one-shot pruning (Frantar & Alistarh, 2023; Sun et al., 2023), global pruning that makes pruning decisions based on all layers (Bai et al., 2024), and domain-specific pruning (Zhang et al., 2024a).
+Recent unstructured pruning includes one-shot pruning (Frantar and Alistarh, 2023; Sun et al., 2023), global pruning that makes pruning decisions based on all layers (Bai et al., 2024), and domain-specific pruning (Zhang et al., 2024a).
 Structured pruning includes gradient-based  (Xia et al., 2024; Ma et al., 2023) and non-gradient-based (Ashkboos et al., 2024) methods.
@@
 Only black-box distillation will work on closed-source models.
+B.5 Module-Wise Compression
+Recent module-wise compression methods (Lu et al., 2024; Yin et al., 2025) are designed to address imbalanced quality among different model components (e.g., layers and weight matrices).
+This research direction could potentially address the problem of over-compressing the important weights, which is also one of our motivations of locating important weights.
+As a way of enriching module-specific compression, we speculate that using non-uniform compression ratios across different modules (e.g., unstructured pruning and mix-precision quantization) might pose challenges for inference speedup.
+A potential solution of being more hardware-efficient is to enforce uniform ratios with protection mechanisms on important weights.
+For example, AWQ protects salient weights identified via activation signals by scaling them up, while still using a uniform compression ratio across modules.
+As the cornerstone of module-wise compression, it is quite valuable to look for novel importance signals for locating salient weights (Liu et al., 2025b).
+For example, our paper decodes the effects of compression on LRMs and validates the usefulness and generalizability of fine-grained mechanistic interpretation for LRMs compression.
+A successful compression method might need to incorporate multiple signals or merge them, as some may benefit from a change in the compression goal (e.g., changing the domain or the compression ratio).
 Table 5: Dataset statistics of selected reasoning benchmarks.
-Size Answer Type Metric Knowledge
+Size Answer Type Metric Knowledge Required?
 AIME 2024 30 Integer Accuracy False
@@
 Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models during inference: maximum generation length is set to $32768$, temperature is set to $0.6$, and top-p value is set to $0.95$.
+Within the 2.51-bit LRM, the embedding matrices, output/head, and normalization/router layers are kept at higher bit widths, while almost everything else is quantized aggressively.
+Specifically, the embedding matrix is in 4-bit and the final language model head is in 6-bit, while all MoE (Mixture of Experts) routing matrices and layer-norm weights are left in full precision (32-bit).
+The majority (around 88%) of weights – namely the bulk of the MoE weights – are quantized down to 2.51-bit.
+We do not find precise details of the calibration dataset used for 2.51-bit quantization.
+However, we see that Unsloth plans to use more than 1.5 million tokens for future quantized models, so we can make a reasonable prediction that the dynamic quantization uses more calibration data than other quantization methods we benchmark in our paper.
+We adopt the same calibration data for all other quantization.
+The dynamically quantized models work on MoE architecture such as R1, while other quantization methods (e.g., AWQ and GPTAQ) are designed for non-MoE with less weight count.
+Therefore, due to different compression targets, we are not able to compare dynamic quantization and other methods under identical calibration.
+In Table 1, we aim to report various methods as comprehensive as possible in order to study the effects of compression on reasoning performance.
 We use AutoAWQ [^7] https://github.com/casper-hansen/AutoAWQ as the AWQ implementation for inference due to its speed advantage (vLLM support), while the original AWQ code [^8] https://github.com/mit-han-lab/llm-awq is used to generate the pseudo-quantized R1 distilled Llama-8B for our analysis in Section 5.
@@
 Appendix F Additional Visualization of Weight Importance and Importance Shift
-All additional figures in Appendix are thoroughly discussed in the main content.
 Figure 4 shows the weight importance of DeepSeek-R1-Distill-Qwen-7B across four heatmaps, each corresponding to a specific target reasoning behavior.
@@
 Figure 7: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit GPTQ quantized variant.
-Appendix G Test-time Compute
+Table 6: Statistics of ratings under 2 temperature values. ”CI” represents 95% confidence interval.
+Behavior Avg (temp=1) Avg (temp=0.1) Spearman $\rho$ CI
+Backtracking 0.9 0.94 0.892 [0.474, 1.000]
+Uncertainty Estimation 0.78 0.75 0.803 [0.416, 1.000]
+Example Testing 0.90 0.87 0.862 [0.645, 1.000]
+Adding Knowledge 0.86 0.96 0.486 [-0.218, 1.000]
+Appendix G Annotation Robustness
+It is important to find a validated approach to automatically perform annotation, as a large number of $s_{i}^{c}$ would help the computation of importance scores.
+We did not change the behavior definitions and instruction prompts.
+The most salient adjustable parameter is the temperature, and we used the default temperature value (1) in our annotation pipeline.
+We study whether the change of temperature would overturn our annotation results.
+We choose a low temperature value (0.1) and see whether more deterministic outputs would affect our annotation results significantly.
+Setting temperature to 0.1, we first use GPT-4o with the exact same behavior definitions and instruction prompts to collect annotation.
+Then, we perform human validation to judge the output quality and stability when a lower temperature is set.
+To score a model output instance on a specific behavior, we instruct the evaluator to go through all the behavior tokens that are tagged and compute the percentage of correctly tagged behavior.
+For example, suppose GPT-4o tags 20 “backtracking” token sequences in an output instance and the evaluator finds 18 of them closely following the definition of “backtracking”, the score of this instance is 0.9.
+There are cases when an instance ends up with no tagged behavior.
+Since we encourage a conservative annotation style (the value of $s_{i}^{c}$ is large enough for computation such as 1000, so it is fine if GPT-4o misses some token sequences of the target behavior), we assign the score of 1 (maximum score) to those instances ending up with 0 tagged behavior.
+For each behavior, the evaluator samples 10 instances out of 120 (with replacement) and scores them.
+We present the average score of 10 sampled instances under 2 temperatures and the Spearman $\rho$ with 95% confidence interval in Table 6.
+As shown, the average rating scores of both temperatures are high across 4 behaviors, which indicates reliable annotation results done by GPT-4o.
+Comparing the results of both temperatures side-by-side, we observe that a temperature of 1 tends to produce more nested and unclosed token sequences, which is expected due to higher randomness.
+This does not affect our interpretation analysis, since we exclude those nested or unclosed sequences.
+As for Spearman $\rho$, we see that the two rating scores on all 4 reasoning behaviors exhibit at least a moderate and often strong correlation, which demonstrates stable rank ordering.
+For “adding knowledge” behavior, the 95% confidence interval is wide, indicating uncertainty due to the small sample size of 10.
+Given high average rating scores, we conclude using the default temperature value does not significantly affect the correctness of annotation.
+Therefore, the robustness of our findings is validated.
+Figure 8: Change of $\mathbf{RI}_{m\ell}^{c}$ (increases only) from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant.
+Figure 9: Change of $\mathbf{RI}_{m\ell}^{c}$ (net change) from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant.
+Appendix H Justification of Visualization Choice
+As elaborated in Section 2.3, we track changes in relative importance from a more reasoning-capable model to a less reasoning-capable model (rather than the other way around).
+Then, we mention to only consider decreases to show our interest in tracking the cases when the reasoning capability of a weight matrix is diminished.
+In other words, when an LRM is compressed via suboptimal pruning or quantization, salient weight components are more likely to show decreased rather than increased relative importance, because suboptimal compression tends to preserve or harm, but not improve, the reasoning capabilities of important components.
+For compression methods that fully preserve or increase the relative importance of some salient components, tracking these increases adds little value, as it essentially confirms that the methods are performing well.
+It is more valuable to visualize the effects of compression with an emphasis on the weakness of current compression methods, which can inspire future improvement such as our third key finding.
+On the other hand, in order to show all aspects of importance shift, we draw the heatmaps of only considering increases (Figure 8) and those that reflect the net change (Figure 9) from R1 distilled Llama-8B to its 4-bit AWQ variant.
+We see that the net-change heatmaps appear quite similar to those that only consider decreases (Figure 3).
+Specifically, the net-change heatmaps primarily highlight only a few components from the increase-only heatmaps based on Figure 3, and our third key finding still holds (researchers are recommended to prioritize the preservation of final-layer-module importance in order to improve current quantization methods).
+The similarity between Figure 3 and net-change heatmaps showcases a significant shortcoming of current quantization, as most of the net change comes from diminishing capabilities of certain salient modules in Figure 3.
+Figure 10: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its AlphaPruning variant (at 30% sparsity).
+Appendix I Pruning Effect on Weights
+As mentioned in Section 3.1, we require a pruned LRM that remains usable for interpretability analysis, as annotating and interpreting outputs that consist largely of gibberish would be meaningless.
+Therefore, in order to provide more insights on pruning, we turn to AlphaPruning (compared to SparseGPT on the R1 distilled Llama-70B, AlphaPruning achieves higher scores on most benchmarks in Table 1).
+On R1 distilled Llama-8B, we iteratively reduce the 50% sparsity level and ultimately settle on 30%, where much stronger reasoning performance is achieved.
+Although 30% sparsity (1.4x reduction in size) offers much smaller compression ratio than 4-bit quantization (4x reduction in size), its “avg” score of 64.1 is significantly higher than 4-bit GPTQ and GPTAQ in Table 1.
+We then select AlphaPruning at 30% sparsity on the R1-distilled Llama-8B as the pruned LRM for our interpretability analysis.
+We show the corresponding heatmaps in Figure 10.
+We observe that Figure 10 appears very similar to 4-bit AWQ on R1 distilled Llama-8B (Figure 3), 4-bit GPTQ on R1 distilled Llama-8B (Figure 7), and 4-bit AWQ on R1 distilled Qwen-7B (Figure 6).
+Therefore, our third key finding can be nicely generalized to AlphaPruning as well: AlphaPruning overly compresses the final-layer modules and MLP gate projections.
+This additional experiment shows the effects of pruning and the generalizability of our third key finding on pruning methods.
+Figure 11: $\mathbf{I}_{m\ell}^{c}$ of DistilGPT-OSS-qwen3-4B.
+Figure 12: Change of $\mathbf{RI}_{m\ell}^{c}$ from DistilGPT-OSS-qwen3-4B to its 4-bit GPTQ variant.
+Appendix J Generalizability of Key Findings on Non-R1 Families
+We are excited to share that our key findings generalize well to Pinkstack/DistilGPT-OSS-qwen3-4B (a distilled model from both GPT-OSS-20B and GPT-OSS-120B via SFT).
+In order to systematically investigate weight importance and quantization effect, we interpret this distilled model and its 4-bit GPTQ quantized variant.
+We first show their one-pass benchmarking scores in Table 7.
+It is clear to see that MuSiQue scores are not affected by quantization, further implying that preserving weight count with lower precisions is a reasonable strategy for LRMs’ knowledge memorization (first key finding).
+Table 7: Performance of DistilGPT-OSS-qwen3-4B and its 4-bit GPTQ variant.
+Model AIME 2024 FOLIO Temporal Avg MuSiQue
+DistilGPT-OSS-qwen3-4B 36.7 82.3 94.4 71.1 (4.0, 6.73)
+4-bit GPTQ quantized DistilGPT-OSS-qwen3-4B 26.7 76.4 86.4 63.2 (4.0, 6.92)
+Then, we look at our second key finding.
+After running our fine-grained interpretation analysis on DistilGPT-OSS-qwen3-4B, we present its heatmaps in Figure 11.
+We see that 36_up (up_projection in the final layer) is either the most important or the second most important component across four heatmaps, which demonstrates the generalizability of our second key finding.
+Moreover, whenever 36_up is ranked second on a heatmap, 36_gate is always ranked first.
+As both up_projection and gate_projection in the final layer are important to R1 distilled models in Figure 2 (upper part) and 4, we see a high similarity of weight importance between R1 distilled models and this GPT-OSS distilled model, which indicates the possibility of finding common important components across different LRMs.
+On the other hand, it is worth noting that different LRMs might end up with slightly different heatmaps.
+For example, last-layer k_projection is highlighted in Figure 4 but appears to be less important for the GPT-OSS distilled model.
+This shows that our interpretation can produce customized analysis of weight importance.
+Our second key finding aims to locate the import weight component that is shared by different LRMs (e.g., different model sizes and classes).
+Finally, for our third key finding, we demonstrate its generalizability through the 4-bit GPTQ quantized DistilGPT-OSS-qwen3-4B.
+The heatmaps for visualizing decrease of relative importance are shown in Figure 12.
+Most salient decreases (outliers) come from the last-layer components across all four heatmaps, along with a few salient decreases in the second last layer.
+Out of a small number of outliers in a heatmap, gate_projection also has at least 2 outliers on the backtracking and uncertainty estimation capabilities.
+Therefore, even when we switch to a GPT-OSS distilled model, GPTQ still overly compresses the final-layer modules and MLP gate projections.
+Figure 13: $\mathbf{I}_{m\ell}^{c}$ of Llama-3.1-8B.
+Appendix K Importance Scores of Base Model
+Complementing the lower part (change of relative importance from R1 distilled Llama-8B to its base model before distillation) of Figure 2, we present the importance scores of the base model (Llama-3.1-8B) in Figure 13.
+As we can see, the heatmaps of the base model look very different from those of the R1-distilled model (upper part of Figure 2).
+There are many differences.
+For example, compared to uncertainty estimation behavior, other reasoning behaviors do not present a clear outlier of importance, which indicates more evenly distributed causal relationships across weight matrices on backtracking, example testing, and adding knowledge.
+As for uncertainty estimation, the last-layer o_projection has the highest importance, whereas the up_projection in the final layer is not even ranked among the top three components of that layer.
+Therefore, we could reach a completely different conclusion from key finding 2 if we were to interpret the base model rather than the LRMs.
+The commonality between the two parts of Figure 2, together with the difference between the base model and the lower part of Figure 2, demonstrates Takeaway 4.3 (“important weights of the R1 distilled models are mainly the result of the distillation effect”).
+Table 8: Performance of naive RAG and closed-book settings on MuSiQue.
+Setting LRM MuSiQue
+Closed-book R1-Distill-Llama-8B (0.0, 4.43)
+RAG R1-Distill-Llama-8B (2.0, 14.52)
+Closed-book R1-Distill-Qwen-7B (0.0, 3.57)
+RAG R1-Distill-Qwen-7B (2.0, 13.54)
+Appendix L Additional RAG Pipeline
+Section 3.3 concentrates on the significantly lower MuSiQue scores as we move from R1 to different compressed variants.
+We hypothesize that parametric knowledge is more sensitive to parameter count, because pruned R1 distilled Llama-70B in Table 2 collapses on MuSiQue even earlier than AIME 2024.
+Other notable phenomena include the slightly stronger reasoning performance of R1 distilled Qwen-32B than Llama-70B, but with collapsed MuSiQue scores at even 0% sparsity.
+In order to further validate our key finding 1, we implement a RAG pipeline across two R1 distilled models with different sizes.
+Since all MuSiQue scores in Table 1 follow the closed-book setting, we provide results from naive RAG to decouple reasoning from knowledge memorization.
+Specifically, we use BAAI/bge-base-en-v1.5 as the embedding model and retrieve the top 10 chunks whose embeddings are closest to each question in MuSiQue.
+We run this RAG pipeline using R1-Distill-Llama-8B and R1-Distill-Qwen-7B to handle questions as shown in Table 8.
+We observe that scores get improved significantly by retrieving relevant knowledge on both LRMs, which demonstrates the lack of memorized knowledge.
+We would like to highlight that RAG with R1-Distill-Qwen-7B even surpasses closed-book R1-Distill-Qwen-32B on F1.
+This provides strong evidence that knowledge memorization is largely compromised when compressing from 32B to 7B (a substantial reduction in parameter count), whereas multihop reasoning capability is not clearly affected by this reduction.
+Therefore, taken together with the discussion in Section 3.3, our key finding 1 is validated.
+Appendix M Additional Diagnosis
+We performed another diagnosis by running AWQ, checking if there is alignment between our interpretation results and AWQ.
+Since a salient weight channel typically gets scaled up more significantly than others, its scaling factor essentially serves as an importance measure that is based on activations.
+Removing its normalization on scaling factors that is done component by component, we extract the scaling factors from 4-bit AWQ on R1 distilled Llama-8B.
+We notice that the last-layer up projection gets assigned to the second largest scaling factor across all up projections by excluding the scaling factor of 1 (scaling factor of 1 means no scaling), so we see a strong alignment between AWQ and our key finding 2.
+The incorporation of the gradient signal in attribution patching increases our importance score of the last-layer up projection from the second largest to the largest.
+Table 9: Performance of selectively quantizing 32_up or 1_up on AIME 2024.
+Selective Quantization 32_up 1_up
+4-bit (R1-Distill-Llama-8B) 20.0 43.3
+5-bit (R1-Distill-Llama-8B) 26.7 30.0
+3-bit (R1-Distill-Qwen-7B) 16.7 46.7
+Appendix N Additional Selective Quantization
+First of all, in Table 3, both $20.0$ and $6.7$ are pretty low scores on AIME 2024 given that only a single component is quantized, since the original model reaches 42.2 over three runs in Table 1.
+We suspect a greater variability on low-score output, so we take a closer look at the outputs of quantizing 32_up and 1_up.
+We find that the output of only quantizing 32_up contains significantly more gibberish than 1_up, which indicates the low output quality of only quantizing 32_up.
+For example, the total output size of only quantizing 32_up is $10.9$ MB while the total output size of only quantizing 1_up is $7.5$ MB.
+Taking these signals into consideration, we additionally demonstrate the lower output quality of only quantizing 32_up on AIME 2024 by performing selective 4-bit and 5-bit quantization on 32_up and 1_up.
+To show generalizability, we also perform selective 3-bit quantization on both components over R1-Distill-Qwen-7B.
+The resulting AIME 2024 scores are shown in Table 9.
+We see that only quantizing 32_up yields lower scores than 1_up on all three cases.
+Considering the lower average scores of only quantization 32_up in Table 3, we think the exception in Table 3 (quantizing only 1_up causes the largest drop on AIME 2024) is due to random perturbation of declined capabilities on AIME 2024.
+Appendix O Test-time Compute
 We study the behavior of R1 and some of its compressed variants by measuring their test-time compute.
@@
 R1-Distill-Qwen Distillation & 30% sparse 100.0 0.0 8.9 87.5 75.0 11.0 (30.0, 36.6) (5.0, 11.2) 13.4
-Appendix H Case Study
+Appendix P Case Study
 As discussed in Section 3.2, pruned models collapse on all benchmarks at certain sparsity levels.